A Complete Bibliography of Publications by, and about, Leslie Fox

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

13 August 2025 Version 1.01

Abstract

This bibliography records publications of Leslie Fox.

Title word cross-reference

(< 40) [Jen59]. 0 [Hof59]. 0 ≤ x < ∞ [Shi73]. 1 000 000 [SL58]. 90 [Hof59]. A [Alt78, Jel76]. Ax = λBx [LM80]. C_{n}(x) [Ano52b]. f(x, y) = ∫_{0}^{∞} exp(−k)(J_{0}(kx) cosh(ky) − 1)cosech(k) dk [F^{+}45b]. f_{λ} [Huc56]. h^{2} [Den64, Osb60]. h^{4} [Den64]. K(1/b_{n}) [Fie78]. LR [DT71]. M [Per65]. n [Rut55]. π [FH75]. S_{n}(x) [Ano52b]. x = 10 [Ano55b, Ano55a]. x = 2 [Ano55a]. x = 2.5 [Ano55b]. x′ = f(t, x)

[Ste71]. y′′ = ϕ(x, y, y′) [Sal57]. Y_{1}(mx)[xJ_{1}(x) − BJ_{0}(x)] − J_{1}(mx)[xY_{1}(x) − BY_{0}(x)] = 0. [LZ56].

-extrapolation [Den64, Osb60]. -stability [Jel76]. -stable [Alt78]. -ter [Rut55]. -th [Rut55].

0 [McC85]. 0-387-90820-X [McC85].

11 [Bue66]. 12th [Jac77]. 13-15 [Cra87]. 15th [Jac77]. 1958 [Lan59]. 1961 [Fox62b]. 1962 [Tod64]. 1965 [Wal66]. 1966 [Gun67]. 1967 [DH69]. 1970 [YFM71]. 1976 [Jac77]. 1985 [Ano85]. 1986 [Fox87b]. 1987 [Cra87]. 1992 [Ano92].

421 [Kuk72a].

63s [Bur67].

Abel [And73]. Abscissas [DR56, Rab60]. Academic [Fox74a]. Accelerating [Per67]. account [Fox48b]. Accuracy [Pai80, BM59, Fox61c, Ste53]. ACE [Rob56]. ACM [Cra87, Kuk72a]. Addison [P.64, Tod64]. Addison-Wesley [P.64, Tod64]. address [Fox85b, Fox65f]. adjoint [Wit58]. adjustment [KS56]. Advances [Fox66a, Bur67, Gun67, McG67]. aerodynamic [Huc56]. aerodynamical [F^{+}45a]. After [Fox90b]. After-Dinner [Fox90b]. Algebra [Bak66, DH69, Fox64, Fox65c, Fox73b, Fox79b]. Algebraic [FHW48a, FHW48b, Alt57, FHV50, Gol72, Tsa75]. Algol [Fox68a]. Algorithm [Kuk72a, Pai80, Bre80, Cha74, DT71, FH68, Hol74, Yam68]. Algorithmic [Con65]. algorithms [Faz73, Fox83, PB74a, PB74b, Per67, TN81]. almost [Kee71]. almost-triangular [Kee71]. alternative [Şch78a]. analogy [dGAFMS42, dGAFMS45]. Analyse [Fox68a]. Analysis [Con65, DG64, Fox51, FM51, FP68, Fox75a, Fox78, Fox93, Hil56, Jac77, Lie68, Sin64, Wal66, DT71, FS45, Fox72, Fox76b, Fox83, Fox84, Fox85a, Fox87a, Fox90a, Ker87, UY56, Cle68, Fox58a]. analytic [TM70]. analytical [Bre77]. angular [Spe53]. Answers [Fox71a]. Appendix [FR62]. application [dGAFMS42, dGAFMS45, dGAR62, dGA62b, And73, And68, Ric11, Ros58, Spe53]. Applications [Wal66, DG64, KK78]. applied [dGAFMS42, dGAFS43, dGAFMS45, dGAFS45, dGA62a, CFG^{+}42, CFG^{+}45, Fox42, FS45, Fox46]. Approach [Con65, dGA62b, May64, Ned70, OS81]. approché [CLM59]. approchée [Bla54, Bla56]. Approximate [Iwa70, Kun77, Bla54, Bla56, CLM59, Kul73, Ric11, Vej60, ZL80]. approximating [Boo55, You56]. Approximation [Hay70, Lan59, Fox65a, Fox65d, Osb64a, YFM71]. Approximationen [Buc56]. Approximations [FHM67, Buc56, DS80, Mil59, Row55, Scr73]. April [Jac77, Lan59]. arch [Fox65e]. Arguments [FM51, Fox54d, Fox60c, L.61, Toc55]. arithmetical [Ric11]. Art [Jac77]. Aspects [DH69]. Associated [Ano60, dGAR62]. August [Fox62b]. Automatic [Fox51, Fox60a, FM51]. aux [Her58].

B [Cle68, CFG^{+}42, CFG^{+}45]. Backward [Fox78, Lem76]. balancing [Gra71]. band [Kee71]. Based [Zar55, Fox62b, Wal66]. Bateman [Fox57b]. beams [FS42]. beginnings [Fox85a]. behind [UY56]. Bemerkungen [Rut55]. Bennett [Fox57b, McC85]. Bernoulli [Bau58]. Bessel [Toc55, Ano52a, Ano60, Fox54d, SA57]. best [Fox75b]. biharmonic [Fox47a, RIY61, FS45]. Binomial [Toc55]. Birmingham [Wal66]. Bivariate [SK58]. Boltzmann [Ken76]. Boltzmann-equation [Ken76]. Book [Bak66, BBB00, Bur67, Cle68, Fox61a, Fox68a, Fox74a, Fox76a, Gun67, Hen59, L.61, McC85, McG67, P.64, Tho59, Toc55, Tod59, Tod64, YFM71]. boundaries [Vis57]. Boundary [FM57, Fox57a, FS69, Fox90c, Hen59, Sel58, Tho59, dGAR62, BMO61, BM59, CO68, Fal68, Fox44, Fox47a, Fox50a, FS73, Fox74b, Fox77a, Fox80, FV80, Fox81, GN66, Has61, Hol74, Kul73, Mun81, Nic60, Per65, Tre62, Vej60]. Boundary-value [FM57, BMO61, Fox77a, Hol74, Tre62, Vej60]. Bounds [FHM67, Fie78, Tre62]. box [FS42]. Bratley [McC85]. Bueckner [GV66].

C [Goo65, Toc55, dGAFMS42, dGAFMS45]. Calcul [Kun58, CLM59]. calculated [UY56]. Calculation [Spe53, Bre77, CLM59, Kun58]. calculations [Lad71, PB74a, PB74b, Rob56]. Call [Ano85]. Canonical [Ort74]. Card [FHW48b, KS56]. Cartesian [SK58]. case [SV53]. centre [CC61]. certain [Fai78, GM65, Kee71, Wil59]. Chairman [Fox85b]. characteristic [BR55]. characteristics [Lis60]. charged [Spe53]. Chebyshev [Fox76a, Ano52b, Chi56, Cle57, Ell60, Fox62a, Fox65a, Fox65b, FP68, Kun77, MH77, Sch78b, Cle68]. Checking [Fox60a]. Christoph [YFM71]. class [Fox67, HW77, Mun81, Per65]. cm [L.61]. cm. [P.64, Tod59]. Co [Tod64]. Coefficients [Sal59, SST60, Fox58b, Hor56, Toc55]. Combination [Bre77]. Commemorative [Fox78]. Comments [Fox68b, GV66, Fox61c, FW66]. common [PB74a]. Company [P.64]. comparatively [Rai69]. complete [Gau58]. Complex [Kuk72a, Kuk72b, Sal55]. Computation [AE71, CH90, Cra87, Fox46, Fox48a, FA51, Fox66a, Fox71a, FH74, Hou55, FS41, Fox60a, Fox61c, Hit74, Bur67, McG67]. Computations [Fox74a, Gun67]. Computer [Fox69a]. Computers [Fox83, Fox59b, Fox76b, SA57]. Computing [And73, Bae69, Fox51, FM51, FGM^{+}57, Fox61b, FR62, FM68, FGM^{+}70, Nas90]. Conditioned [Gau75, Men56]. conditions [dGAR62, BM59, Fal68, Fox44, Fox47a, Fox50a, Has61]. conducted [DH69]. Conference [Cra87, Jac77]. conformal [dGA62a]. conjugate [TN81]. connected [Faz73]. consistent [NF69]. constant [PK72]. Constants [Tod75a, Tod75b, Tod00, FH74]. Construction [Fox56, Fox59a, Mil49, Tod59]. Constructive [DH69]. contact [dGA62b]. continued [Fie78]. Control [Kuk72a, Kuk72b, Alt78]. convenience [Fox61c]. Convergence [DK73, Per67]. convergent [BM36]. Convexity [YFM71]. convolution [Jon61]. coordinates [RIY61]. CORDIC [Hit74]. corner [Wig69]. correction [Per65]. corrections [CR80]. Correspondence [FY63]. corresponding [AE71]. Corrigendum [Tod75a]. Cosecants [Rey54]. Cosines [Ano55a]. Course [Lie68]. criteria [BQ76]. curved [UY56, Vis57]. Cylinder [Fox60c, L.61].

D

[Bae69, GV66, dGAFS43, dGAFS45, FH68]. D. [Bue66]. dam [Ric11]. dams [Fox65e]. dans [Ili66]. Data [Hay70, And73]. Deferred [CR80, May64]. Definite [FH70]. Degrees [Hof59, War57]. Dennis [Goo65]. d’équations [Bla54, Bla56]. derivative [Fox50a]. derivatives [HW60, RMF77, Row55, Str60, de 55]. Descending [Ano55b]. Descent [Fis53]. Determination [Fox65b, FS46]. determine [dGA62b, Hol74]. developing [Wit58]. diagonalization [Yam68]. Difference [Fox53, Fox61a, Row55, dGA62b, BMO61, BM59, Cha76, Cla66, Fox61c, Fox77a, Fox81, Her58, Hol74, Ken76, Mun81, Osb64a, Per65]. differences [Fox79a, Ric11, Her58]. Differential [Fis53, Fox57a, Fox57b, Fox61a, Fox62b, FM87, Fox90c, Hen59, Mil53, P.64, Pol64, Sel58, Tho59, Tod64, dGAR62, dGA62b, Alt78, Bla54, CO68, cC70, Cla66, Cle57, Den60, Fal68, Fox47b, FG49b, FG49a, Fox50a, Fox50b, Fox54a, FM57, Fox60b, Fox62a, Fox63, Fox68b, FS69, FMOT71, Fox71b, FV80, FM81, Gea65, GPP80, Has61, KWC76, Lis60, May64, MH77, NOT76, Ola77, OS81, Osb64a, RIY61, Ric11, Rut55, Shi73, Ste53, Usm72, Wit58, You56, ZL80, de 55]. Differentialgleichungen [Rut55]. differentiation [PK72]. différentielles [Bla54]. difficulty [SV53]. digit [Gau58]. Digital [Hou55, Fox59b]. dimensional [dGAFMS42, dGAFMS45, dGAR62, FS41, Iri68]. Dimensions [YFM71, Lad71].

Dinner [Fox90b]. d’intégration [Bla54]. Dirichlet [Wig69]. discretization [Per67]. Discussions [Fox51]. distributions [dGAFS43, dGAFS45, dGA62b, Spe53]. divisor [PB74a]. Do [Fox71a]. d’ordre [Her58]. double [FHM73]. Dover [Fox57b]. Dr [Gau58]. Dr. [Fox87c, Gau58]. Dr.-Ing [Gau58]. Dr.-Ing. [Gau58].

each [Orl74a]. Early [Fox87a, Fox90a]. Edinburgh [Gun67]. Edited [Bur67, McG67, Tod64, YFM71]. Editorial [DW^{+}93]. education [Fox84, Fox68a]. effect [BM59, Ste71]. Effectiveness [Pai80]. Efficient [PB74a, PB74b, Fai78, GPP80]. Eigenproblem [Pai80]. eigenvalue [Den64, Fox60b, FHM73, Fox77b, GM65, Osb60, Osb64b, SV53]. Eigenvalues [FHM67, Gyr74, AE71, Nic67, Wit58]. eigenvectors [AE71]. elastic [dGAFS43, dGAFS45, CFG^{+}42, CFG^{+}45, FS45, FA51, SV53]. Elasticity [Bue66, GV63]. electro [dGAFMS42, dGAFMS45]. electro-magnetic [dGAFMS42, dGAFMS45]. electrodynamics [Kul73]. Elementary [Con65, Zar55, Men56]. Elements [Sin64]. eleven [Kay55]. Elliptic [FHM67, Fox50a, FS69, Fox71b, Fox77a, Fox79a, Fox81, Mun81, Nic67]. end [CR80]. engineering [dGAFMS42, dGAFS43, dGAFMS45, dGAFS45, CFG^{+}42, CFG^{+}45, FS45]. Engineers [FM68, Bae69]. England [Tod59, Wal66]. Entries [Ano85]. Epilogue [Fox90b]. Equation [Zar55, And73, Bar52, BR55, Fal68, FMOT71, Ken76, KWC76, LZ56, Scr73, Ste72, Vis57, You56]. Equations [Fis53, FHW48a, FHW48b, Fox50c, Fox53, Fox54b, Fox57a, Fox62b, FM87, Fox90c, Mil53, dGAR62, dGA62b, Alt78, Alt57, BFMW64, Bla54, Bla56, Bol72, Cha76, CO68, cC70, Cla66, Cle57, Den60, DS80, Ell60, FHV50, F^{+}45a, Fox47b, FG49b, FG49a, Fox50a, Fox50b, FG53, Fox54a, Fox54c, FM57, Fox60b, Fox62a, Fox63, Fox68b, FS69, Fox71b, FV80, FM81, Gea65, GPP80, Gol72, Has61, Her58, HW77, Jon61, KK78, Lis60, May64, Men56, MH77, NOT76, Ola77, OS81, Osb64a, RIY61, Ric11, RS56, Rut55, Shi73, Ste53, Usm72, Wil59, Wit58, ZL80, dHW74, de 55, P.64, Her58, Fox57b, Fox61a, Hen59, Pol64, Sel58, Tho59, Tod64]. Equilibration [Lem76]. equipment [KS56]. erreur [Ili66]. Error [Fie78, Fox78, Hou55, KG68, Kuk72a, Kuk72b, TM70, Kay55, PK72, Sal55, Sch78b]. Errors [YFM71, Kul73]. Escalator [Fox52]. Estimating [Kul73]. estimation [TM70]. Evaluation [Fox69a, RMF77, Ili66]. Everett [Chi56, Fox58b]. exceeding [FS44, SL58]. exceptional [SV53]. exemplified [SV53]. Exercises [Fox65c, Bak66]. Existence [BQ76]. expansion [Heg75]. expansions [CC61, Fox65b, Mun81, Nic60, Sch78b, Spe53]. experiments [Fox60b, Fox71b, FV80, You56]. explicit [de 55]. Exponential [Ano55b, FM51, Mil60, Den60, Hit74, Led61]. expressible [SL58]. Extension [Nic67, FS45]. extensions [Ste68]. extrapolated [DK73]. extrapolation [Bre80, Den64, Osb60]. Extremal [Mil59].

F [Bae69]. F.R.S. [Fox87c]. factor [NF69]. factorization [FH68, Ker87]. factorizations [BQ76]. Falsi [FS73]. Fast [Cha74]. Fermi [Buc56]. Fermi-Funktion [Buc56]. FFT [Cha74]. field [UY56]. Filon [Cha74]. finies [CLM59]. Finite [Cla66, Fox53, Fox77a, Fox79a, Fox81, YFM71, dGA62b, BMO61, BM59, CLM59, Fox61c, Mun81, Osb64a, Ral78, Ric11, Fox61a]. Finite-Difference [Fox53, Cla66, Fox77a, Fox81, dGA62b, BMO61, Fox61c, Osb64a]. finite-rank [Ral78]. first [BFMW64, Fox54a, HW77, Kay55, de 55]. first-order [Fox54a]. Five [Gau58]. Five-digit [Gau58]. flat [CFG^{+}42, CFG^{+}45, FS45]. flexure [FS45]. flow [FS44, Huc56, Jae56, Jen59]. flux [UY56]. fonctions [Kun58]. Foreward [FW75]. form [MH57, PB74a]. Forming [Orl74a]. formula [Chi56]. formulae [Wit58]. Formulas [Sal56a, Sal56b, Bla54, Bol72, CR80, HW60, Iwa70, PK72, Str60]. formules [Bla54]. Forsythe [Fox61a]. four [SL58]. Fourier [Cha74, Lad71]. fourth [KWC76]. fourth-order [KWC76]. Fox [Ano92, Bae69, Bak66, Bur67, Cle68, Gun67, L.61, McC85, McG67, P.64, Tho59, Toc55, Tod59, Tod64, Ano85, Cle92, DW^{+}93, Hen59, Ker87, MW99, McK18, Nic93, Nic67, OW18, Pol64, Sel58]. fractions [Fie78]. frameworks [CFG^{+}42, CFG^{+}45]. Frankfurt [Gun67]. Fredholm [BFMW64, Bol72, Ral78]. free [FS73, Fox74b, dGAFMS42, dGAFMS45]. free-boundary [FS73, Fox74b]. freedom [War57]. French [Bla54, Bla56, CLM59, Her58, Kun58]. Function [Fox69a, Kuk72a, Kuk72b, Buc56, F^{+}45b, Heg75, Iwa70, Kay55, Ral78, Sal55]. functional [FMOT71]. Functions [Ano52a, Ano60, Fox54d, Fox60c, Hay70, L.61, Toc55, Boo55, Hit74, Huc56, Kun58, Led61, Orl74b, RI70, Sch78b, SA57, TM70]. Fundamental [DH69]. Fünfstellige [Gau58]. Funktion [Buc56]. Further [FW66, FH75].

G [Fox61a, Fox74a]. Galerkin [DS80, HW77]. Gamma [Kuk72a, Kuk72b]. gas [FS44]. Gaussian [DR56]. General [Fox66b, Bre80]. Generalized [Mil60, NF69]. generated [Gyr74]. Generation [Hou55, SA57]. German [Buc56, Gau58, Rut55, Tre62]. Get [Fox71a]. gewöhnlicher [Rut55]. Gobbin [Gau58]. good [Fox83]. gradient [dGAR62, TN81, Vis57]. Graeffe [Bau58]. Green [Ral78]. Grid [SK58]. Griffin [Bue66]. Guide [McC85].

H [Fox57b, Fox87c, Gau58, YFM71].

Hankel [CLM59]. Hardy [Fox87b]. Hayes [Ker87]. heat [Jae56, Scr73]. Held [Fox62b, Jac77]. helium [Bar52]. helping [FH75]. Henrici [Nic67]. Her [L.61, Tod59]. Herausgegeben [Gau58, Gun67]. Hermitian [Bos75]. heuristic [Ste71]. high [DR56, Rab60, SA57]. higher [Fox84, Her58]. Higher-order [Her58]. Hill [Fox58a, Zar55]. History [Cra87, Nas90]. hooks [FS46]. HSNC’87 [Cra87]. Hybrid [Gea65]. hydrogen [CC61]. Hyperbolic [Ano55a, DS80, Hit74, Lis60].

IBM [DH69]. II [Ano52a, PB74b]. iii [L.61, Ano60]. ill [Men56]. Imbedding [KK78]. implementation [Fai78]. Implicit [dHW74, FM81]. improvement [ZL80]. improvements [Fox47b]. incorrect [Ned70]. indefinite [KG68]. Ing [Gau58]. Ing. [Gau58]. inital [GN66]. initial [FM57, Gea65]. initial-value [FM57]. instability [SV53]. Institute [Wal66]. Integer [Fox54d, Toc55]. integers [SL58]. Integral [FM51, BFMW64, Bol72, Ell60, F^{+}45a, FG53, HW77, Jon61, KK78, Ste72, dHW74]. intégrales [CLM59]. Integrals [Mil60, CLM59, Kay55, KG68]. integrand [HW60, Str60]. Integrands [FH70, Fox67]. Integration [Fox57b, FH70, RI70, Rut55, SST60, Bla54, Bla56, Den60, Fal68, FG49b, Fox54a, Fox67, Fox68b, Kun77, Led61, Sal57, Shi73, TM70, Usm72, de 55, Bla56]. integro [Cha76, NOT76]. integro-difference [Cha76]. integro-differential [NOT76]. Interpolation [SK58, Sal59, Chi56, Fox58b, Iri68, Sal56a, Sal57]. interval [FV80]. Intervals [Fox53]. Introduction [Bak66, Fox64, Fox65c, Fox70, Fox73b, Fox79b, Hil56, Wal66, Fox74a]. Introductory [Fox75a]. inverse [Fox77b, Heg75, Sal56a]. Inversion [Fox50c, FH51, Fox54b, Fox54c, KS56]. inverting [Gol72, MH57, RS56]. involve [Fox50a]. involving [HW60, Ric11, Str60]. ion [CC61]. ISBN [McC85]. Iteration [Bau58]. Iterative [Bar52, Fox73a, Nic60, You71, BQ76, GPP80, Tsa75, ZL80]. ix [P.64, Ric11].

J [Bue66, Fox76a, Fox87c, Toc55]. James [Fox87b]. Jersey [Cra87]. Jim [Fox78, Fox90b]. John [Fox61a, Fox76a]. Josef [YFM71]. June [DH69]. jury [War57].

Kaiser [Fox58a]. kernels [Ral78]. kind [BFMW64, HW77, Kun77, Orl74b, Shi81, dHW74]. Kingdom [Fox87a, Fox90a]. known [Gyr74]. Kronecker [Ste68]. Kunz [Fox58a]. Kutta [Shi81, dHW74].

L [Ano85, Bak66, Bur67, Cle68, Gun67, Hen59, McC85, McG67, Pol64, Tho59, Toc55, Tod64]. L. [Bae69, Ker87, L.61, P.64, Sel58, Tod59]. Lab [Tod59]. Laboratory [DH69, L.61]. Lanczos [Ort74, Pai80, Yam68]. Large [FM51, Fox53, Fox54d, Fox60c, L.61, Toc55, You71, BR55, MH57, Wil59]. latent [Fox52]. Laurent [Heg75]. layer [Fox46, FA51]. learn [Fox83]. Least [Fox65d, Fox65a, Lot56]. Lees [Şch78a]. Legendre [Kun58, Spe53]. Lemniscate [Tod75a, Tod75b, Tod00, FH74]. Leslie [Cle92, McK18, Nic93, Ano92, DW^{+}93, MW99, OW18]. limit [May64]. linéaires [Ili66]. Linear [Bak66, FHW48a, FHW48b, Fox50c, Fox54b, Fox64, Fox65c, Fox73b, Fox79b, YFM71, You71, Alt57, CO68, Cle57, Fai78, Fis53, FHV50, FG53, Fox54c, FS69, Fox71b, Gol72, Has61, Iri68, Men56, PB74a, PB74b, Per65, RS56, Ste68, Tsa75, Wit58]. lines [KWC76]. Linus [McC85]. Liouville [GM65]. liquid [Lad71]. load [FA51]. loading [Fox46]. Lobatto [Rab60]. location [Bre68]. logarithmic [Gau58, Hit74]. logarithmische [Gau58]. London [Fox61a, Gun67, L.61, Tod59]. Lösungen [Tre62]. lot [Fox83]. low [Jen59]. Ltd [Fox61a]. LX [BM36].

Machinery [Fox51, FM51]. Machines [Fox61b]. made [Orl74b]. Madison [Lan59]. magnetic [dGAFMS42, dGAFMS45]. Majesty [L.61, Tod59]. Making [FM51]. many [War57]. masonry [Ric11]. Mass [P.64]. Math [Tod59]. Mathematical [Fox56, Fox69a, Fox69b, Fox85b, L.61, Mil49, Orl74b, Tod59]. Mathematics [Wal66]. Mathieu [BR55, Zar55]. Matrices [Fox50c, FH51, Fox54b, Gau75, Bos75, Bre68, Faz73, Fox54c, Gol72, Gyr74, Kee71, Lem76, Men56, MH57, PB74a, RS56]. Matrix [Bau58, Faz73, Fox74a, Gra71, BQ76, Gyr74, Heg75, KS56, TN81, Wil59]. May [Cra87]. Mayers [Bae69]. McGraw [Fox58a]. McGraw-Hill [Fox58a]. mean [Iwa70]. Meaningless [Fox71a]. Mechanical [FR62]. membranes [dGAFMS42, dGAFMS45]. Merbach [YFM71]. mesh [BM59]. Method [Fis53, And68, cC70, DK73, Fal68, FS73, Hit74, HW77, Iri68, Ker87, KS56, Kul73, KWC76, Lis60, MH57, NOT76, Ola77, Ort74, OS81, Osb64a, Osb64b, Per65, RS56, UY56, Usm72, Vis57, Wig69, Wit58, de 55]. Methods [dGA54, Bae69, FHW48b, Fox50c, FH51, Fox51, Fox54b, FGM^{+}57, Fox61a, FM68, FGM^{+}70, dGAFMS42, dGAFS43, dGAFMS45, dGAFS45, dGAR62, dGA62a, Alt78, BQ76, BMO61, Bre77, Bro73, Cha76, CFG^{+}42, CFG^{+}45, Fai78, FS41, Fox42, Fox44, FS45, FS46, Fox47b, Fox48b, FG49b, FG49a, Fox52, Fox59a, Fox62a, Fox65e, Fox75b, Fox77a, Fox80, FV80, Fox81, Gea65, Gol72, Jel76, KK78, LM80, Rai69, Ros58, Shi81, Squ60, Wil59, dHW74]. Miller [Toc55]. Milne [Fox57b]. Min [Fox68a]. Minimal [PK72, PB74b]. Minimax [Fox59a]. minimization [LM80]. Mixed [Fox47a, Fox44]. modelling [Fox85b]. Modern [Bau58, FGM^{+}57, FGM^{+}70]. Modification [Gol72, Bro73]. modifications [Ral78]. modified [Chi56]. modulated [RI70]. molecular [CC61]. Moler [Nic67]. Müller [YFM71]. Müller-Merbach [YFM71]. multi [Iri68, Usm72]. multi-dimensional [Iri68]. multi-step [Usm72]. multiderivative [Jel76]. multidimensional [Bos75]. multiple [AE71, Spe53]. multiplication [KS56]. Multiplikations [Fox73a]. multistep [Jel76]. multivariate [RMF77]. My [McK18].

N.D [PK72]. National [L.61, Tod59]. Nationale [Fox68a]. Natural [Hof59]. near [dGA62b, Kun77]. negativity [Bos75]. network [Row55]. Neumann [Wig69]. Newton [Bro73, NOT76]. no [Bue66]. noisy [And73]. Non [Bos75, Bur67, Fis53, Fox66a, Gun67, McG67, FG53, Has61, Per65]. Non-linear [Fis53, Has61, Per65]. Non-negativity [Bos75]. Non-Numerical [Bur67, Fox66a, Gun67, McG67]. non-singular [FG53]. nonlinear [cC70, NOT76, OS81]. nonsymmetric [TN81]. normal [dGAR62, RI70, Vis57]. normal-gradient [dGAR62]. Note [FS41, Goo65, Jel76, Lot56, And68, Boo55, Fai78, Fox54a]. Notes [FHV50, FHW48a, Fox51, FM51]. nozzle [FS44]. nuclear [Rob56]. Number [Fox73a]. numbers [Jen59, SL58]. Numeric [Cra87]. Numerical [BFMW64, Bak66, Bue66, Bur67, Cha76, Cle68, Con65, CH90, DG64, Fis53, Fox51, FM51, Fox57a, Fox62b, Fox64, Fox65c, Fox65e, Fox66a, FP68, Fox73b, Fox75a, Fox76b, Fox79b, Fox80, Fox84, FM87, Fox90c, Fox93, GV63, Gun67, Hay70, Hen59, Hil56, Hol74, Jac77, Jae56, Lad71, Lan59, Led61, Lie68, McG67, Mil53, Sal57, Sel58, Shi73, Sin64, Ste72, Tho59, Wal66, Zar55, BM36, BMO61, Bol72, Bre77, CO68, Cle57, Den60, Ell60, FG49b, Fox50a, Fox50b, FG53, Fox54a, FM57, Fox60b, Fox68b, Fox72, Fox75b, FM81, Fox83, Fox85a, Fox85b, Fox87a, Fox90a, GPP80, Jon61, Ken76, Lem76, Lis60, MH77, Orl74a, Orl74b, OS81, PK72, Rut55, Ste53, Ste71, TM70, Usm72, de 55, P.64, Fox57b, Fox58a, Pol64, Tod64]. Numérique [Fox68a]. numerischen [Rut55].

Obituary [Cle92, Fox87c, MW99, Nic93]. occur [Huc56]. occurring [F^{+}45a]. oceanic [Ros58]. October [Fox87b]. off [Wig69, YFM71]. Office [L.61, Tod59]. one [Alt78, Lad71]. one-step [Alt78]. operational [OS81]. operations [Orl74a]. Operators [FHM67, Hol74, Nic67, Ste68]. Optimally [Gau75]. Optimization [YFM71]. optimize [Lem76]. optimum [NF69]. order [DR56, DS80, Fox54a, Her58, KWC76, Rab60, Rut55, de 55]. ordering [NF69]. Orders [Fox54d, Toc55]. Ordinary [Fox57a, Fox62b, FM87, Fox90c, Hen59, P.64, Pol64, Sel58, Tho59, Tod64, Alt78, cC70, Den60, Fal68, Fox47b, FG49b, FG49a, Fox50b, FM57, Fox60b, Fox62a, FV80, FM81, Gea65, May64, Ola77, Osb64a, Rut55, Ste53, Usm72, Wit58, ZL80]. Ordnung [Rut55]. organized [Wal66]. orthogonal [Fox65d]. oscillating [Huc56]. oscillations [dGAFMS42, dGAFMS45]. Osculatory [SK58, Sal59, SST60, Sal56a, Sal57]. Other [Fox60c, L.61, Ral78]. over-relaxation [NF69]. overtone [And68]. Oxford [Bak66, Cle68, Gun67, Tho59, Fox62b, OW18]. Oxford-London-Edinburgh-New [Gun67].

P [Toc55, L.61, P.64, Tod59]. Padé [Fai78]. pages [Cle68, Fox61a, Fox68a, Fox74a, Fox76a, Tod64]. Paper [Bue66, GV66, Goo65, Nic67]. Paperback [L.61]. papers [Cra87]. Parabolic [Fox60c, L.61, Bla56, Fai78]. parabolique [Bla56]. parameter [BR55]. Paris [Gun67]. Parker [Cle68]. Part [Ano52a, Ano60]. Partial [Fox61a, Fox62b, Fox63, P.64, Pol64, Tod64, dGAR62, BQ76, Cla66, Fox47b, Fox71b, GPP80, Hor56, KWC76, Lis60, Row55, Sal56b]. particles [Spe53]. partition [cC70]. partitioned [Bre68, Gyr74]. Paul [McC85]. peaked [Spe53]. Pegasus [Wil59]. Pergamon [Bur67, Gun67, McG67]. Perturbed [Vej60]. Phase [Rob56]. Physical [L.61, Tod59, Fox69b, Ric11]. Pi [BBB00]. Piecewise [cC70]. Pilot [Rob56]. Place [Hof59]. Plane [Bue66, GV63, CFG^{+}42, CFG^{+}45, Fox42, Fox44, Fox47a]. plates [FS45]. plating [CFG^{+}42, CFG^{+}45]. Point [Hen59, Sel58, Tho59, BM59, CO68, Fal68, Fox57a, FV80, Fox90c, GN66, Has61]. points [Kun77]. Poisson [Vis57]. polar [RIY61]. pole [Kun77]. Polynomial [FH68, Scr73, Ker87, Orl74a, PB74a, Spe53]. Polynomials [Cle68, FP68, Ano52b, Boo55, cC70, Ell60, Fox65d, Fox69b, MH77, Ort74, RMF77, Fox76a]. positive [BM36]. possessing [Den60]. potential [Fox44]. pour [CLM59, Her58]. pp [McC85]. pp. [Bak66, Fox57b, Fox58a, McG67, Tho59]. Practical [Fox50c, FH51, Fox54b, Fox54c]. Preconditioned [TN81]. preconditioning [GPP80]. Presentations [Pol64]. presented [Cra87]. Presidential [Fox65f]. Press [Bak66, Cle68, Fox74a, Gun67, McG67, Tho59]. Princeton [Cra87]. Prize [Ano85]. problem [F^{+}45a, FHM73, Fox76b, Fox77b, Fox79c, Wig69]. Problems [Bue66, Fox57a, Fox90c, GV63, Hen59, Sel58, Tho59, dGAFMS42, dGAFS43, dGAFMS45, dGAFS45, BMO61, BM59, CO68, CFG^{+}42, CFG^{+}45, Den64, Fai78, Fox42, Fox44, FS45, Fox47a, FM57, Fox60b, FS73, Fox74b, Fox77a, Fox79a, Fox80, FV80, Fox81, Gea65, GM65, GN66, Hol74, KS56, Kul73, Mun81, Ned70, Nic60, Osb60, Osb64b, Per65, Ric11, SV53, Tre62, Vej60, War57]. procedures [Bar52, Fox68a]. Proceedings [Jac77, Cra87, DH69, Lan59]. Processes [Bau58, Fox73a, Fox61c]. product [Bol72, Ste68]. Professor [Cle92, McK18, Nic93]. Programming [Bur67, Fox66a, Gun67, McG67, YFM71]. projection [GN66]. proof [Şch78a]. proper [Fox65f]. properties [DK73, Fox65b, Men56]. pseudo [Shi81]. pseudo-Runge [Shi81]. Publications [Fox57b, Fox58a, Pol64, Sel58]. Published [Fox68a, Gau58]. Publishing [P.64, Tod64]. pulsation [And68]. Punched [FHW48b, KS56].

Q [FH68]. Quadrature

[HW60, Bol72, Cha74, Rab60, Str60]. quadratures [DR56]. Quasi [Bro73]. Quasi-Newton [Bro73]. quotient [LM80].

R [Fox61a, Goo65, GV66]. R. [Bue66]. radial [Jae56]. Randwertaufgaben [Tre62]. rank [Ral78]. rates [KG68]. Rayleigh [LM80]. Rayleigh-quotient [LM80]. Reading [P.64]. Real [Zar55]. realization [PB74b]. recollections [McK18]. recurrence [Fox65f, HT73]. Récurrences [Her58]. reduce [Ste71]. Reducing [Fox73a]. Reduction [GN66]. refinement [Tsa75]. reflections [RIY61]. Regula [FS73]. Regula-Falsi [FS73]. related [Sch78b]. relation [HT73]. relations [Fox65f]. Relaxation [dGAFMS42, dGAFS43, dGAFMS45, dGAFS45, dGA54, dGA62a, CFG^{+}42, CFG^{+}45, Fox42, FS45, Fox51, Squ60, dGAR62, DK73, FS41, Fox44, FS46, Fox47b, Fox48b, FG49a, NF69, RIY61, Ros58, Vis57]. relaxational [Fox47a]. Reliable [CH90]. remainder [Nic60]. Remark [Bue66, GV66]. Remarks [Rut55]. repeated [Kay55]. repetitions [Her58]. Report [Gun67, You56]. representation [dGA62b, Cla66]. Research [DH69, Fox61b]. résolution [Ili66]. Resolvent [Ral78]. result [Şch78a]. resulting [Fox46]. Review [Bak66, Bur67, Cle68, Fox61a, Fox68a, Fox74a, Fox76a, Hen59, L.61, McC85, McG67, P.64, Tho59, Toc55, Tod59, Tod64, Zar55]. Reviews [Bae69, YFM71]. revolution [dGAFS43, dGAFS45]. Reynolds [Jen59]. Ritz [And68]. Rivlin [Fox76a]. road [Fox48a]. Romberg [Fox67]. root [Bre68, Hit74, HT73]. root-location [Bre68]. roots [Fox52, LZ56]. rotational [UY56]. Round [YFM71, Jen59]. Round-off [YFM71]. Runge [Shi81, dHW74]. Rüschlikon [DH69].

S [Bue66, Fox58a, GV66, Gun67]. S. [Goo65]. S.E.R. [DK73]. S14 [Kuk72a, Kuk72b]. satisfy [dGAR62]. scattering [Rob56, Spe53]. scheme [GPP80, Ken76]. schemes [Mun81]. School [Fox62b]. Schrage [McC85]. Schranken [Tre62]. Scientific [Cra87, Fox59b, Fox71a, Nas90, DG64, Fox61c]. Scientists [Bae69, FM68]. second [DS80, Kun77, dHW74, de 55]. second-order [DS80]. self [Wit58]. self-adjoint [Wit58]. sensitivity [Lot56]. Sentiments [Fox90b]. September [Fox62b, Fox87b]. Series [GM65, BM36, Cle57, Hor56, Sal56b]. sets [Wil59]. shifted [DT71]. shock [UY56]. Short [Fox54d, Toc55, Fox48b]. SIAM [Bue66]. simple [Fox48a, Kun77, Wit58]. Simulation [McC85]. Simultaneous [FHW48a, FHW48b, LM80, FHV50]. Sine [Hof59]. Sines [Ano55a]. Single [CC61]. Single-centre [CC61]. Singular [FH70, FG53, Fox67]. singularities [Fox68b, FS69, Fox71b, Fox74b, Fox79a]. singularity [Wig69]. size [Alt78, BM59]. sliding [dGA62b]. slowly [BM36]. Smith [PB74a]. Snippet [Fox78]. solids [dGAFS43, dGAFS45]. Solution [Bue66, Fis53, Fox44, F^{+}45a, FHW48a, FHW48b, Fox50c, Fox54b, Fox57a, Fox62b, FM87, Fox90c, GV63, Hen59, Mil53, P.64, Pol64, Sel58, Tho59, Tod64, Vis57, Wil59, You71, Alt57, BFMW64, BMO61, Bol72, CO68, Cle57, Ell60, FHV50, Fox47b, FG49a, Fox50a, Fox50b, FG53, Fox54c, FM57, Fox68b, FM81, GM65, Has61, Jon61, Ken76, Kul73, KWC76, Lis60, MH77, Nic60, Ola77, OS81, Osb64b, Ric11, Scr73, Ste72, Ste71, Vej60, War57, You56, ZL80]. Solutions [Zar55, BM59, Den60, GPP80, Lot56, Ste53, Tre62]. solving [Fox76b, Gol72, RS56]. Some [Fox47b, FG49b, Fox60b, Fox61c, Fox71b, FV80, Men56, RIY61, SV53, ZL80, Fox90b]. Sons [Fox61a]. sound [FS44]. Source [BBB00]. space [GN66]. spacings [Row55]. special [MH57]. specified [Vis57]. spectra [Orl74b]. speed [FS44, SA57]. sphere [Jen59]. Springer [McC85]. Springer-Verlag [McC85]. square [Hit74, HT73]. Squares [Rey54, Fox65a, Fox65d, Lot56]. stability [CFG^{+}42, CFG^{+}45, Jel76, Lem76]. stable [Alt78, Rai69]. State [Jac77, Lad71]. Stationery [L.61, Tod59]. Steepest [Fis53]. Stefan [Fox79c]. stellar [And68]. Step [Fox51, Alt78, Usm72]. Step-by-Step [Fox51]. step-size [Alt78]. Stewart [Fox74a]. stiff [Alt78]. still [Fox83]. Stoer [YFM71]. strain [Fox42, Fox47a]. stress [FS41, Fox42, Fox47a, dGAFS43, dGAFS45]. stresses [Fox46, FS46, Fox48a, FA51, Ric11]. stressing [FS42]. structure [Fox48a]. Sturm [GM65]. subject [Fal68]. subtract [Wig69]. successive [DK73, NF69]. such [Jon61]. suggested [dGA62b]. sum [SL58]. summary [Mil59]. summation [BM36, Hor56, Sal56b]. Summer [Fox62b]. sums [Led61]. superconvergence [DS80]. supérieur [Her58]. supersonic [Huc56]. surface [Fox46]. Survey [Fox66b]. Switzerland [DH69]. Symmetric [Pai80]. symmetrical [Faz73]. symmetrically [Fox46]. Symposium [Lan59, DH69, Wal66]. system [dGAFMS42, dGAFMS45, Fox46, FA51, PB74a, PB74b]. systèmes [Ili66]. Systems [You71, Alt78, Fal68, FS41, Gol72, Jon61, Tsa75].

Table [Ano55b, Ano55a, BR55, FM51, Fox54d, Hor56, Rey54, SL58, Toc55, Fox59a, Kay55, Orl74a, Toc55]. Table-Making [FM51]. Tables [Ano52b, CLM59, Fox56, Fox58b, Fox60c, Hof59, L.61, Mil60, SK58, Sal59, SST60, Str60, Tod59, Gau58, KG68, Mil49, Orl74b, CLM59, L.61, Tod59]. Tabulation [F^{+}45b, Huc56]. Tafeln [Gau58]. Talbot [YFM71]. Tangent [Hof59]. tapered [FS42]. tasks [Faz73]. tau [Ort74, OS81]. Taylor [Nic60]. Tchebyschev [Ola77]. Teaching [Fox61b, Fox72]. techniques [FM57]. temperature [dGA62b, Jae56]. temperature-distributions [dGA62b]. Ten [Hof59]. ter [Rut55]. terms [BM36, Nic60]. test [Bos75]. tetrahedral [SL58]. tetrahedron [Iwa70]. th [Rut55]. their [FS46, RMF77, Vej60]. Theodore [Fox76a]. Theorem [DH69]. theorems [Bre68]. theory [And68, Huc56, Ned70, YFM71]. third [Shi81]. Thomas [Buc56, Buc56]. three [dGAR62, FA51, Lad71]. three-dimensional [dGAR62]. three-layer [FA51]. tides [Ros58]. traffic [Fox48a]. Transform [Cha74]. transformations [dGA62a]. transforms [Lad71]. transverse [dGAFMS42, dGAFMS45]. treatment [Fox47a]. tri [Yam68]. tri-diagonalization [Yam68]. triangular [Kee71, Tsa75]. tridiagonalization [Rai69]. Trigonometric [Hof59, Boo55, Gau58]. trigonometrische [Gau58]. Two [Fox57a, Fox74b, Fox90c, Hen59, Sel58, Tho59, dGAFMS42, dGAFMS45, BM59, CO68, Fal68, FS41, Fox46, FV80, GN66, Has61, Lad71, Orl74a]. two-dimensional [dGAFMS42, dGAFMS45, FS41]. two-layer [Fox46]. Two-Point [Hen59, Sel58, Tho59, Fox57a, Fox90c, BM59, CO68, Fal68, FV80, GN66, Has61]. Type [Bau58, Bla56, Bol72, Den60].

uncentered [CR80]. uneven [Row55]. Uniform [Mun81, Fox46]. United [Fox87a, Fox90a]. University [Bak66, Cle68, Jac77, Tho59, FW66]. unstable [Fal68]. upper [Fox46]. Use [Fox53, Fox56, FS41, Fox47b, Fox65f, Str60, Tod59]. uses [Fox59b]. using [Bol72, CR80, Ell60, Kun77, KWC76, Led61, Nic60, Sal57, Wil59].

v [Tod59]. value

[BMO61, CO68, FM57, Fox77a, Fox80, FV80, Fox81, Gea65, GN66, Hol74, Iwa70, Kul73, Mun81, Nic60, Orl74a, Per65, Tre62, Vej60]. Values [Ano60, BR55, Jae56, Orl74a]. Vandermonde [Gau75]. Varga [Bue66, GV66]. Variable [Zar55]. variational [Kul73]. vector [GN66, RS56]. velocities [FS44]. Verlag [McC85]. vibrations [dGAFMS42, dGAFMS45]. VII [dGAFMS42, dGAFS43, dGAFMS45, dGAFS45, CFG^{+}42, CFG^{+}45, FS45]. viii [Gun67, McG67]. Viscous [Jen59]. vollständige [Gau58]. Volterra [NOT76, Ral78, Ste72, dHW74]. Volume [L.61].

W [Fox57b, Fox61a, Fox74a]. Wasow [Fox61a]. wave [Bar52, UY56]. Weber [L.61, Fox60c]. weights [DR56, Rab60]. Wesley [P.64, Tod64]. which [Huc56]. Wiley [Fox61a, Fox76a]. Wilkinson [Fox87b, Fox87c, Fox78, Fox90b]. wings [Huc56]. without [de 55]. Witzgall [YFM71]. work [FW66]. wrapping [Ste71].

X [McC85]. xi [Bak66, Tho59]. xv [Fox58a]. xviv [McC85].

York [Bak66, Fox57b, Fox58a, Fox61a, Gun67, Jac77, McC85, McG67, Tho59]. York-Paris-Frankfurt [Gun67].

zeros [Sal55, Ano60]. zur [Rut55]. Zürich [DH69]. Zürich-Rüschlikon [DH69].

References

Andrew:1971:CEC

[AE71] A. L. Andrew and G. C. Elton. Computation of eigenvectors corresponding to multiple eigenvalues. Bulletin of the Australian Mathematical Society, 4:419–422, 1971. CO-DEN ALNBAB. ISSN 0004-9727 (print), 1755-1633 (electronic).

Altman:1957:SLA

[Alt57] M. Altman. On the solution of linear algebraic equations. Bulletin de l’Académie polonaise des sciences. Classe troisième, 5:93–97, IX, 1957. CODEN BPSMA6. ISSN 0554-5897.

Alt:1978:SOS

[Alt78] René Alt. A-stable one-step methods with step-size control for stiff systems of ordinary differential equations. Journal of Computational and Applied Mathematics, 4(1):29–35, 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Andrew:1968:NRM

[And68] A. L. Andrew. A note on the Ritz method with an application to overtone stellar pulsation theory. Journal of the Australian Mathematical Society, 8 (2):275–286, May 1968. CO-DEN JAUMAX. ISSN 0004-9735 (print), 2059-9234 (electronic).

Anderssen:1973:CND

[And73] R. S. Anderssen. Computing with noisy data with an application to Abel’s equation. In Error, approximation, and accuracy (Proc. Sem., Austral. Nat. Univ., Canberra, 1972), pages 61–80. University of Queensland Press, St. Lucia, Brisbane, QLD, Australia, 1973.

Anonymous:1952:BFP

[Ano52a] Anonymous. Bessel Functions, Part II, volume 10 of British Association Mathematical Tables. Cambridge University Press, Cambridge, UK, 1952. ???? pp.

Anonymous:1952:TCP

[Ano52b]  Anonymous, editor.  Ta-

bles of Chebyshev polynomials, S_{n}(x) and C_{n}(x), volume 9 of National Bureau of Standards Applied Mathematics Series. United States Government Printing Office, Washington, DC, USA, 1952. xxx + 161 pp. LCCN ???? Foreword by J. C. P. Miller. Introduction by Cornelius Lanczos.

Anonymous:1955:THS

[Ano55a] Anonymous. Table of Hyperbolic Sines and Cosines, x = 2 to x = 10, volume 45 of National Bureau of Standards Applied Mathematics Series. United States Government Printing Office, Washington, DC, USA, 1955. v + 81 pp.

Anonymous:1955:TDE

[Ano55b]  Anonymous.  Table of the

Descending Exponential, x = 2.5 to x = 10, volume 46 of National Bureau of Standards Applied Mathematics Series. United States Government Printing Office, Washington, DC, USA, 1955. v + 76 pp.

Anonymous:1960:BFP

[Ano60] Anonymous. Bessel Functions.

Part III: Zeros and Associated Values, volume 7 of Royal Society Mathematical Tables. Cambridge University Press, Cambridge, UK, 1960. lx + 79 pp. Prepared under the direction of the Bessel Functions Panel of the Mathematical Tables Committee.

Anonymous:1985:FPC

[Ano85] Anonymous. 1985 L. Fox prize:

Call for entries. IMA Journal of Numerical Analysis, 5 (1):1, January 1985. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Anonymous:1992:LF

[Ano92]  Anonymous.  Leslie Fox:

1918–1992. Oxford University booklet., 1992. ii + 77

pp. URL https://people.
maths.ox.ac.uk/wathen/fox/
Leslie_Fox_booklet_1994.

pdf. Includes brief curriculum vitae (completed 3 days before Leslie Fox’s death), and a bibliography of his publications. Memories of Leslie Fox by C. W. Clenshaw, pp. 7–10.

Baer:1969:RCM

[Bae69] R. M. Baer. Reviews: Computing Methods for Scientists and Engineers, by L. Fox and D. F. Mayers. American Mathematical Monthly, 76(10): 1163, December 1969. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.
jstor.org/stable/2317225.

Baker:1966:BRI

[Bak66] Christopher T. H. Baker. Book review: An Introduction to Numerical Linear Algebra with Exercises. By L. Fox. Oxford University Press, New York, 1965. xi + 327 pp. SIAM Review, 8 (1):125–127, January 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bartlett:1952:IPH

[Bar52] James H. Bartlett. Iterative procedures and the helium wave equation. Physical Review (2), 88(3):525–526, November 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Bauer:1958:MMI

[Bau58] F. L. Bauer. On modern matrix iteration processes of Bernoulli and Graeffe type. Journal of the ACM, 5(3):246–257, July 1958. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735x (electronic).

Berggren:2000:PSB

[BBB00] Lennart Berggren, Jonathan Borwein, and Peter Borwein, editors. Pi: a Source Book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 0-387-98946-3 (hardcover). xx + 736 pp. LCCN QA484 .P5 2000.

Baker:1964:NSF

[BFMW64] C. T. H. Baker, L. Fox, D. F. Mayers, and K. Wright. Numerical solution of Fredholm integral equations of first kind. The Computer Journal, 7(2):141–148, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/7/2/141.full.pdf+
html;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_02/070141.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/
141.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/
142.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_07/Issue_02/tiff/
143.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/
144.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/
145.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/
146.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/
147.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/
148.tif.

Blanc:1954:FIA

[Bla54] Charles Blanc. Sur les formules d’intégration approchée d’équations différentielles. (French) [On the formulas for approximate integration of differential equations]. Archiv der Mathematik, 5:301–308, 1954. CO-DEN ACVMAL. ISSN 0003-889X (print), 1420-8938 (electronic).

Blanc:1956:IAE

[Bla56] Charles Blanc. Sur l’intégration approchée d’équations du type parabolique. (French) [On the approximate integration of parabolic type equations]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 7(2):146–152, March 1956. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Bickley:1936:LNS

[BM36] W. G. Bickley and J. C. P.

Miller. LX. The numerical summation of slowly convergent series of positive terms. Philosophical Magazine, 22(149): 754–767, November 1936. CO-DEN PHMAA4. ISSN 0031-

8086.  URL https://www.
tandfonline.com/doi/abs/
10.1080/14786443608561726.

Borwein:1959:EBC

[BM59] David Borwein and Andrew R.

Mitchell. The effect of boundary conditions and mesh size on the accuracy of finite difference solutions of two-point boundary problems. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 10(3):221– 232, May 1959. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Bickley:1961:FDM

[BMO61] W. G. Bickley, S. Michaelson, and M. R. Osborne. On finitedifference methods for the numerical solution of boundaryvalue problems. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 262(1309): 219–236, July 1961. CODEN PRLAAZ. ISSN 0962-8444 (print), 2053-9177 (electronic).

Boland:1972:NSF

[Bol72] W. Robert Boland. The numerical solution of Fredholm integral equations using product type quadrature formulas. BIT (Nordisk tidskrift for informationsbehandling), 12(1): 5–16, March 1972. CODEN BITTEL, NBITAB. ISSN 0901-246X.

Booth:1955:NAP

[Boo55] A. D. Booth. A note on approximating polynomials for trigonometric functions. Mathematical Tables and Other Aids to Computation, 9(49):21–23, January 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bose:1975:NNT

[Bos75] N. K. Bose. Non-negativity test of multidimensional Hermitian matrices. Journal of The Franklin Institute, 299(6): 453–456, June 1975. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

Beauwens:1976:ECP

[BQ76] Robert Beauwens and Lena Quenon. Existence criteria for partial matrix factorizations in iterative methods. SIAM Journal on Numerical Analysis, 13 (4):615–643, September 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Blanch:1955:TCV

[BR55] Gertrude Blanch and Ida Rhodes. Table of characteristic values of Mathieu’s equation for large values of the parameter. Journal of the Washington Academy of Sciences, 45 (6):166–196, June 1955. CO-DEN JWASA3. ISSN 0043-

0439.  URL https://www.
jstor.org/stable/24533649.

Brenner:1968:NRL

[Bre68] J. L. Brenner. New rootlocation theorems for partitioned matrices. SIAM Journal on Applied Mathematics, 16(5): 889–898, September 1968. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712x (electronic).

Bremi:1977:CAN

[Bre77] P. Bremi. Combination of analytical and numerical calculation methods. In Numerical analysis (Proc. Colloq., Lausanne, 1976), volume 37 of Internat. Ser. Numer. Math., pages 195–208. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1977. ISBN 3-7643-0939-3.

Brezinski:1980:GEA

[Bre80] C. Brezinski. A general extrapolation algorithm. Numerische Mathematik, 35(2): 175–187, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Broyden:1973:QNM

[Bro73] C. G. Broyden. Quasi-Newton, or modification methods. In Numerical solution of systems of nonlinear algebraic equations (NSF-CBMS Regional Conf., Univ. Pittsburgh, Pittsburgh, Pa., 1972), pages 241–280. Academic Press, New York, USA, 1973.

Buchdahl:1956:ATF

[Buc56] H. A. Buchdahl. Über Approximationen der Thomas–Fermi-Funktion. (German) [On approximations of the Thomas– Fermi function]. Annalen der Physik 6 (Berlin, Germany), 17:238–241, 1956. CODEN AN-PYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Bueckner:1966:RPD

[Bue66] Hans F. Bueckner. Remark on a paper by D. S. Griffin and R. S. Varga “Numerical Solution of Plane Elasticity Problems” [J. SIAM 11 (1963), no. 4, 1046–1062]. SIAM Journal on Applied Mathematics, 14(2):414, March 1966. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [GV63] and comment [GV66].

Burstall:1967:BRA

[Bur67] R. M. Burstall. Book review: Advances in Programming and Non-Numerical Computation. Edited by L. Fox. 63s. 1966. (Pergamon). Mathematical Gazette, 51(377):277– 278, October 1967. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL http://www.jstor.org/
stable/3613292.

Cohen:1961:SCE

[CC61] M. Cohen and C. A. Coulson.

Single-centre expansions for the hydrogen molecular ion. Mathematical Proceedings of the Cambridge Philosophical Society, 57 (1):96–106, January 1961. CO-DEN PCPSA4. ISSN 0008-1981. Appendix by L. Fox.

Chu:1970:PPP

[cC70] Shih chi Chu. Piecewise polynomials and the partition method for nonlinear ordinary differential equations. Journal of Engineering Mathematics, 4:65– 76, 1970. CODEN JLEMAU. ISSN 0022-0833 (print), 1573-2703 (electronic).

Christopherson:1942:RMA

[CFG^{+}42] D. G. Christopherson, L. Fox, J. R. Green, F. S. Shaw, and R. V. Southwell. Relaxation methods applied to engineering problems, VII B, the elastic stability of plane frameworks and of flat plating. Philosophical Transactions of the Royal Society, Series C, 1(??):57–83, ???? 1942.

Christopherson:1945:RMA

[CFG^{+}45] D. G. Christopherson, L. Fox, J. R. Green, F. S. Shaw, and R. V. Southwell. Relaxation methods applied to engineering problems. VII B. The elastic stability of plane frameworks and of flat plating. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 239(810):461–487, October 1945. CODEN PTR-MAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Cox:1990:RNC

[CH90] M. G. Cox and S. J. Hammarling, editors. Reliable Numerical Computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$45.00.

Chang:1974:FQF

[Cha74] Gau Zu Chang. On the Filon quadrature and Fast Fourier Transform (FFT) algorithm. Bulletin of the Institute of Mathematics. Academia Sinica, 2(1):21–27, 1974. ISSN 0304-9825 (print), 2304-7917 (electronic).

Chaudhry:1976:NMI

[Cha76] Muhammad Anwar Chaudhry. Numerical methods for integrodifference equations. Memoirs of the Faculty of Science, Kyushu Imperial University. Series A, Mathematics = Kyushu Teikoku Daigaku Rigakubu kiyo, 30(1):15–23, 1976. CODEN MFKAAF. ISSN 0373-6385 (print), 1883-2172 (electronic).

Chisnall:1956:MCE

[Chi56] G. A. Chisnall. A modified Chebyshev–Everett interpolation formula. Mathematical Tables and Other Aids to Computation, 10(54):66–73, April 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Clark:1966:FDR

[Cla66] M. M. Clark. Finite-difference representation of partial differential equations. Buletinul Institutului Politehnic din Iaşi. Serie Nouă, 12/16:31–37, 1966. ISSN 0032-6100.

Clenshaw:1957:NSL

[Cle57] C. W. Clenshaw. The numerical solution of linear differential equations in Chebyshev series. Mathematical Proceedings of the Cambridge Philosophical Society, 53(1):134–149, January 1957. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Clenshaw:1968:BRC

[Cle68] C. W. Clenshaw. Book review: Chebyshev Polynomials in Numerical Analysis, by L. Fox and I. B. Parker, 1969; 205 pages (Oxford University Press). The Computer Journal, 11(3):310, March 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/11/3/310.full.pdf+
html.

Clenshaw:1992:OPL

[Cle92] Charles Clenshaw. Obituary:

Professor Leslie Fox. The Independent, ??(??):??, August 10, 1992. URL https:

                                        //www.independent.co.uk/
                                        news/people/obituary-professor-
                                        leslie-fox-1539620.html.

Chatelet:1959:TPC

[CLM59] F. Châtelet, G. Lansraux, and B. Million. Tables pour le calcul approché des intégrales finies de Hankel. (French) [Tables for the approximate calculation of finite Hankel integrals]. Chiffres: Revue de l’Association française de Calcul, 2:233–237, 1959. ISSN 0245-9922.

Chaves:1968:NST

[CO68] T. Chaves and E. L. Ortiz. On the numerical solution of two-point boundary value problems for linear differential equations. Zeitschrift für Angewandte Mathematik und Mechanik, 48(6):415–418, January 1968. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Conte:1965:ENA

[Con65] S. D. Conte. Elementary Numerical Analysis: an Algorithmic Approach. McGraw-Hill, New York, NY, USA, 1965. x + 278 pp.

Christiansen:1980:DCU

[CR80] J. Christiansen and R. D. Russell. Deferred corrections using uncentered end formulas. Numerische Mathematik, 35(1): 21–33, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Crane:1987:HAC

[Cra87] G. E. Crane, editor. HSNC’87:

ACM Conference on the History of Scientific and Numeric Computation, conference proceedings: papers presented at the Conference, Princeton, New Jersey, May 13-15, 1987. ACM Press, New York, NY 10036, USA, October 1987. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

deVogelaere:1955:MNI

[de 55] Rene de Vogelaere. A method for the numerical integration of differential equations of second order without explicit first derivatives. Journal of Research of the National Bureau of Standards (1934), 54:119–125, 1955. ISSN 0160-1741 (print), 2376-5259 (electronic).

Dennis:1960:NIO

[Den60] S. C. R. Dennis. The numerical integration of ordinary differential equations possessing exponential type solutions. Mathematical Proceedings of the Cambridge Philosophical Society, 56 (3):240–246, July 1960. CO-DEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Dennis:1964:EEP

[Den64] S. C. R. Dennis. h^{2}- and h^{4}-extrapolation in eigenvalue problems. Mathematical Proceedings of the Cambridge Philosophical Society, 60(1):67– 74, January 1964. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Dodes:1964:NAS

[DG64] Irving Allen Dodes and Samuel L. Greitzer. Numerical Analysis. With scientific applications. Hayden Book Co., Rochelle Park, NJ, USA, 1964. ix + 390 pp.

Allen:1954:RM

[dGA54] D. N. de G. Allen. Relaxation Methods. McGraw-Hill, New York, NY, USA, 1954. ix + 257 pp.

Allen:1962:RMA

[dGA62a] D. N. de G. Allen. Relaxation methods applied to conformal transformations. Quarterly Journal of Mechanics and Applied Mathematics, 15(1):35– 42 (1 insert), 1962. CODEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Allen:1962:SAF

[dGA62b] D. N. de G. Allen. A suggested approach to finite-difference representation of differential equations, with an application to determine temperaturedistributions near a sliding contact. Quarterly Journal of Mechanics and Applied Mathematics, 15(1):11–33, 1962. CO-DEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Allen:1942:RMA

[dGAFMS42] D. N. de G. Allen, L. Fox, H. Motz, and R. V. Southwell. Relaxation methods applied to engineering problems. VII C. Free transverse vibrations of membranes, with an application (by analogy) to twodimensional oscillations in an electro-magnetic system. Philosophical Transactions of the Royal Society, Series C, 1(??): 85–97, ???? 1942.

Allen:1945:RMAc

[dGAFMS45] D. N. de G. Allen, L. Fox, H. Motz, and R. V. Southwell. Relaxation methods applied to engineering problems. VII C. Free transverse vibrations of membranes, with an application (by analogy) to twodimensional oscillations in an electro-magnetic system. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 239(810):488–500, October 1945. CODEN PTR-MAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Allen:1943:RMA

[dGAFS43] D. N. de G. Allen, L. Fox, and R. V. Southwell. Relaxation methods applied to engineering problems. VII D. Stress distributions in elastic solids of revolution. Philosophical Transactions of the Royal Society, Series C, 1(??):99–135, ???? 1943.

Allen:1945:RMAd

[dGAFS45] D. N. de G. Allen, L. Fox, and R. V. Southwell. Relaxation methods applied to engineering problems. VII D. Stress distributions in elastic solids of revolution. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 239 (810):501–537, October 1945. CODEN PTRMAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Allen:1962:ARM

[dGAR62] D. N. de G. Allen and B. Robins. The application of relaxation methods to satisfy normal-gradient boundary conditions associated with threedimensional partial differential equations. Quarterly Journal of Mechanics and Applied Mathematics, 15(1):43–51, 1962. CO-DEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Dejon:1969:CAF

[DH69] Bruno Dejon and Peter Henrici, editors. Constructive Aspects of the Fundamental Theorem of Algebra: proceedings of a symposium conducted at the IBM Research Laboratory, Zürich-Rüschlikon, Switzerland, June 5–7, 1967. Wiley, New York, NY, USA, 1969. ISBN 0-471-20300-9. LCCN QA212 .C65.

deHoog:1974:IRK

[dHW74] F. de Hoog and R. Weiss. Implicit Runge–Kutta methods for second kind Volterra integral equations. Numerische Mathematik, 23(3):199–213, June 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

DellaTorre:1973:CPS

[DK73] E. Della Torre and W. Kinsner.

Convergence properties of the successive extrapolated relaxation (S.E.R.) method. Journal of the Institute of Mathematics and its Applications, 12(3):175– 185, October 1973. CODEN JMTAA8. ISSN 0020-2932.

Davis:1956:AWG

[DR56] P. Davis and P. Rabinowitz.

Abscissas and weights for Gaussian quadratures of high order. Journal of Research of the National Bureau of Standards (1934), 56:35–37, 1956. ISSN 0160-1741 (print), 2376-5259 (electronic).

Dougalis:1980:SGA

[DS80] Vassilios A. Dougalis and Steven M. Serbin. On the superconvergence of Galerkin approximations to second-order hyperbolic equations. SIAM Journal on Numerical Analysis, 17(3):431–446, June 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dekker:1971:ASA

[DT71] T. J. Dekker and J. F. Traub. An analysis of the shifted LR algorithm. Numerische Mathematik, 17(3):179–188, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Duff:1993:ELF

[DW^{+}93] I. S. Duff, G. A. Watson, et al. Editorial: Leslie Fox, 1918– 1992. IMA Journal of Numerical Analysis, 13(1):i–ii, January 1993. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Elliott:1960:NSI

[Ell60] David Elliott. The numerical solution of integral equations using Chebyshev polynomials. Journal of the Australian Mathematical Society, 1(3):344–356, August 1960. CODEN JAU-MAX. ISSN 0004-9735 (print), 2059-9234 (electronic).

Fox:1945:SIE

[F^{+}45a] L. Fox et al. Solution of integral equations occurring in an aerodynamical problem. Report SRE/ACS 89, Admiralty Computing Service (S.R.E. Dept.), Bath, UK, 1945. ???? pp.

Fox:1945:TF

[F^{+}45b]  L. Fox et al.  Tabulation

of the function f(x, y) = ∫_{0}^{∞} exp(−k)(J_{0}(kx) cosh(ky) − 1)cosech(k) dk. Report SRE/ACS 47, Admiralty Computing Service (S.R.E. Dept.), Bath, UK, 1945. ???? pp.

Fox:1951:CLS

[FA51] L. Fox and W. E. A. Acum.

Computation of load stresses in a three-layer elastic system. Geotechnique Inst Civ Eng, ?? (??):293–300, ???? 1951.

Fairweather:1978:NEI

[Fai78] Graeme Fairweather. A note on the efficient implementation of certain Padé methods for linear parabolic problems. BIT (Nordisk tidskrift for informationsbehandling), 18(1):106– 109, March 1978. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Falkenberg:1968:MIU

[Fal68] J. C. Falkenberg. A method for integration of unstable systems of ordinary differential equation subject to two-point boundary conditions. BIT (Nordisk tidskrift for informationsbehandling), 8(4):86–103, October 1968. CODEN BIT-TEL, NBITAB. ISSN 0901-246X.

Fazekas:1973:MAT

[Faz73] F. Fazekas. Matrix algorithms for tasks connected with symmetrical matrices. Zeitschrift für Angewandte Mathematik und Mechanik, 53(12):T186– T188, January 1973. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Fox:1949:SRM

[FG49a] L. Fox and E. T. Goodwin. The solution by relaxation methods of ordinary differential equations. Mathematical Proceedings of the Cambridge Philosophical Society, 45(1):50– 68, January 1949. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Fox:1949:SNM

[FG49b] L. Fox and E. T. Goodwin. Some new methods for the numerical integration of ordinary differential equations. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 45(3): 373–388, July 1949. CODEN PCPSA4. ISSN 0008-1981.

Fox:1953:NSN

[FG53] L. Fox and E. T. Goodwin. The numerical solution of nonsingular linear integral equations. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 245(902): 501–534, February 1953. CO-DEN PTRMAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Fox:1957:MCM

[FGM^{+}57] L. Fox, E. T. Goodwin, J. G. L. Michel, F. W. J. Olver, and J. H. Wilkinson. Modern Computing Methods, volume 16 of NPL Notes on Applied Science. Her Majesty’s Stationery Office, London, UK, 1957. vi + 128 pp. LCCN QA297 .T4.

Fox:1970:MCM

[FGM^{+}70] L. Fox, E. T. Goodwin, J. G. L.

Michel, F. W. J. Olver, and J. H. Wilkinson. Modern Computing Methods, volume 16 of NPL Notes on Applied Science. Her Majesty’s Stationery Office, London, UK, second edition, 1970. ISBN 0-11-480021-

9. vii + 169 pp. LCCN QA297 .T4 1970.

Fox:1951:MPM

[FH51] L. Fox and J. G. Hayes. More practical methods for the inversion of matrices. Journal of the Royal Statistical Society. Series B (Methodological), 13 (1):83–91, ???? 1951. CODEN JSTBAJ. ISSN 0035-9246.

URL http://links.jstor.
org/sici?sici=0035-9246(
1951)13:1<83:MPMFTI>2.0.
CO;2-V&origin=MSN;   http:
//www.jstor.org/stable/2983969.

Fox:1968:PFQ

[FH68] L. Fox and Linda Hayes. Polynomial factorization and the Q– D algorithm. Linear Algebra and its Applications, 1(3):445– 463, July 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0024379568900190. See anal-

ysis [Ker87].

Fox:1970:DIS

[FH70] L. Fox and Linda Hayes. On the definite integration of singular integrands. SIAM Review, 12(3):449–457, July 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fox:1974:CLC

[FH74] L. Fox and Linda Hayes. Computation of the lemniscate constants. Bulletin of the Institute of Mathematics and its Applications, 10(??):425–426, ???? 1974. CODEN IMTABW. ISSN 0950-5628.

Fox:1975:FH

[FH75] L. Fox and Linda Hayes. A further helping of π. Mathematical Gazette, 59(407): 38–40, March 1975. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

                                        tronic). URL https://www.
                                        jstor.org/stable/3616808.

Fox:1967:ABE

[FHM67] L. Fox, P. Henrici, and C. B. Moler. Approximations and bounds for eigenvalues of elliptic operators. SIAM Journal on Numerical Analysis, 4 (1):89–102, March 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fox:1973:DEP

[FHM73] L. Fox, Linda Hayes, and D. F. Mayers. The double eigenvalue problem. In J. J. H.

Miller, editor, Topics in numerical analysis (Proceedings of the Royal Irish Academy Conference, University College, Dublin, 1972), pages 93–112. Academic Press, New York, USA, 1973.

Foks:1950:NSS

[FHV50] L. Foks, H. D. Haski, and Dž. H. Vilkinson. Notes on the solution of simultaneous linear algebraic equations. Uspekhi Matematicheskikh Nauk (N.S.), 5(3(37)):60–86, 1950. CODEN UMANA5. ISSN 0042-1316 (print), 2305-2872 (electronic). Russian translation of [FHW48a].

Fox:1948:NSA

[FHW48a] L. Fox, H. D. Huskey, and J. H. Wilkinson. Notes on the solution of algebraic linear simultaneous equations. Quarterly Journal of Mechanics and Applied Mathematics, 1(1):149– 173, January 1948. CODEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Fox:1948:SAL

[FHW48b] L. Fox, H. D. Huskey, and J. H.

Wilkinson. The solution of algebraic linear simultaneous equations by punched card methods. Maths, National Physical Laboratory, Teddington, Middlesex TW11 0LW, UK, 1948.

Field:1978:EBC

[Fie78]  David A. Field.  Error

bounds for continued fractions K(1/b_{n}). Numerische Mathematik, 29(3):261–267, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fischbach:1953:NSN

[Fis53] Joseph W. Fischbach. The Numerical Solution of Nonlinear Differential Equations by the Method of Steepest Descent. Ballistic Research Laboratories, Aberdeen Proving Ground, MD, USA, 1953. 23 pp. Memo. Rep. No. 646.

Fox:1951:NNAb

[FM51] L. Fox and J. C. P. Miller. Notes on numerical analysis — 5: Table-making for large arguments. the exponential integral (in Automatic Computing Machinery). Mathematical Tables and Other Aids to Computation, 5(35):163–166, July 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Fox:1957:BVT

[FM57] L. Fox and A. R. Mitchell. Boundary-value techniques for the numerical solution of initialvalue problems in ordinary differential equations. Quarterly Journal of Mechanics and Applied Mathematics, 10(2):232– 243, 1957. CODEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Fox:1968:CMS

[FM68] L. Fox and D. F. Mayers. Computing Methods for Scientists and Engineers. Monographs on Numerical Analysis. Clarendon Press, Oxford, UK, 1968. xii + 255 pp.

Fox:1981:NSI

[FM81] L. Fox and D. F. Mayers. On the numerical solution of implicit ordinary differential equations. IMA Journal of Numerical Analysis, 1(4):377–401, October 1981. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Fox:1987:NSO

[FM87] L. Fox and D. F. Mayers. Numerical Solution of Ordinary Differential Equations. Chapman and Hall, Ltd., London, UK, 1987. ISBN 0-412-22650-2. xii + 249 pp. LCCN QA372 .F69 1987.

Fox:1971:FDE

[FMOT71] L. Fox, D. F. Mayers, J. R. Ockendon, and A. B. Tayler. On a functional differential equation. Journal of the Institute of Mathematics and its Applications, 8(3):271–307, December 1971. CODEN JMTAA8. ISSN 0020-2932.

Fox:1942:RMA

[Fox42] Leslie Fox. Relaxation methods applied to problems of plane stress and strain. D.Phil. dissertation, University of Oxford, Oxford, UK, 1942. 79

pp.  URL https://www.
proquest.com/pqdtglobal/
docview/2497394794.

Fox:1944:SRM

[Fox44] L. Fox. Solution by relaxation methods of plane potential problems with mixed boundary conditions. Quarterly of Applied Mathematics, 2(3): 251–257, ???? 1944. CODEN QAMAAY. ISSN 0033-569X (print), 1552-4485 (electronic).

Fox:1946:CST

[Fox46] L. Fox. Computation of the stresses in a two-layer system resulting from uniform loading symmetrically applied to the upper surface. Road Research Laboratory Report RN/723, Her Majesty’s Stationery Office, London, UK, 1946.

Fox:1947:MBC

[Fox47a] L. Fox. Mixed boundary conditions in the relaxational treatment of biharmonic problems (plane strain or stress). Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 189(1019):535–543, June 1947. CODEN PRLAAZ. ISSN 0962-8444 (print), 2053-9177 (electronic).

Fox:1947:SIU

[Fox47b] L. Fox. Some improvements in the use of relaxation methods for the solution of ordinary and partial differential equations. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 190(1020):31–59, June 1947. CODEN PRLAAZ. ISSN 0962-8444 (print), 2053-9177 (electronic).

Fox:1948:CTS

[Fox48a] L. Fox. Computation of traffic stresses in a simple road structure. Road Res Lab Tech Paper 9, Her Majesty’s Stationery Office, London, UK, 1948.

Fox:1948:SAR

[Fox48b] L. Fox. A short account of relaxation methods. Quarterly Journal of Mechanics and Applied Mathematics, 1(1):253– 280, January 1948. CODEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Fox:1950:NSE

[Fox50a] L. Fox. The numerical solution of elliptic differential equations when the boundary conditions involve a derivative. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 242(849):345–378, May 1950. CODEN PTRMAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Fox:1950:NSO

[Fox50b] L. Fox. The numerical solution of ordinary differential equations. Proc I.C.M., 1(??):??, ???? 1950.

Fox:1950:PMS

[Fox50c] L. Fox. Practical methods for the solution of linear equations and the inversion of matrices. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):120–136, ???? 1950. CODEN JSTBAJ. ISSN 0035-9246. URL http://

links.jstor.org/sici?sici=
0035-9246(1950)12:1<120:
PMFTSO>2.0.CO;2-W&origin=
MSN;   http://www.jstor.
org/stable/2983838.

Fox:1951:NNAa

[Fox51] L. Fox. Notes on numerical analysis — 4: Relaxation and step-by-step methods (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 5(34):92–95, April 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Fox:1952:EML

[Fox52] L. Fox. Escalator methods for latent roots. Quarterly Journal of Mechanics and Applied Mathematics, 5(2):178– 190, 1952. CODEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Fox:1953:ULI

[Fox53] L. Fox. The use of large intervals in finite-difference equations. Mathematical Tables and Other Aids to Computation, 7 (41):14–18, January 1953. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Fox:1954:NNI

[Fox54a] L. Fox. A note on the numerical integration of first-order differential equations. Quarterly Journal of Mechanics and Applied Mathematics, 7(3):367– 378, 1954. CODEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Fox:1954:PMS

[Fox54b]  L. Fox.  Practical Meth-

ods for the Solution of Linear Equations and the Inversion of Matrices, volume 39 of NBS Applied Mathematics Series. United States Government Printing Office, Washington, DC, USA, 1954. 1–54 pp.

Fox:1954:PSL

[Fox54c] L. Fox. Practical solution of linear equations and inversion of matrices. In Contributions to the solution of systems of linear equations and the determination of eigenvalues, volume 39 of National Bureau of Standards Applied Mathematics Series, pages 1–54. United States Government Printing Office, Washington, DC, USA, 1954.

Fox:1954:STB

[Fox54d] L. Fox. A Short Table for Bessel Functions of Integer Orders and Large Arguments, volume 3 of Royal Society Shorter Mathematical Tables. Cambridge University Press, Cambridge, UK, 1954. 28 pp.

Fox:1956:UCM

[Fox56] L. Fox. The Use and Construction of Mathematical Tables, volume 1 of N.P.L. Mathematical Tables. Her Majesty’s Stationery Office, London, UK, June 1956. iv + 60 pp.

Fox:1957:NST

[Fox57a] L. Fox. The Numerical Solution of Two-point Boundary Problems in Ordinary Differential Equations. Monographs on Numerical Analysis. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1957. xi + 371 pp. Reprinted in [Fox90c].

Fox:1957:RPN

[Fox57b] L. Fox. Recent publications: Numerical Integration of Differential Equations, A. E. Bennett, W. E. Milne, H. Bateman, Dover, New York, 1956, 108 pp. American Mathematical Monthly, 64(7):520, August/September 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fox:1958:RPN

[Fox58a]  L. Fox.  Recent publica-

tions:Numerical Analysis, by Kaiser S. Kunz, McGraw-Hill, New York, 1957. xv + 381 pp. American Mathematical Monthly, 65(6):464–465, June/ July 1958. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972 (electronic).

Fox:1958:TEI

[Fox58b] L. Fox. Tables of Everett interpolation coefficients. Her Majesty’s Stationery Office, London, UK, 1958. 61 pp.

Fox:1959:MMT

[Fox59a] L. Fox. Minimax methods in table construction. In Langer [Lan59], pages 233–244. LCCN QA3 .U45 no. 1. Publication no. 1 of the Mathematics Research Center, U.S. Army, the University of Wisconsin.

Fox:1959:SUD

[Fox59b] L. Fox. Scientific uses of digital computers. Bulletin of the Institute of Physics and the Physical Society, ??(??): 185–188, ???? 1959.

Fox:1960:CAC

[Fox60a] L. Fox. Checking in automatic computation. Inst. Elec. Eng. and British Computer Society, ??(??):14–16, ???? 1960.

Fox:1960:SNE

[Fox60b] L. Fox. Some numerical experiments with eigenvalue problems in ordinary differential equations. In R. E. Langer, editor, Boundary Problems in Differential Equations, pages 243– 255. The University of Wisconsin Press, Madison, WI, USA, 1960.

Fox:1960:TWP

[Fox60c]  L. Fox.  Tables of We-

ber Parabolic Cylinder Functions and Other Functions for Large Arguments, volume 4 of National Physical Laboratory Mathematical Tables. Her Majesty’s Stationery Office, London, UK, 1960. iii + 40 pp. Department of Scientific and Industrial Research.

Fox:1961:BRF

[Fox61a] L. Fox. Book review: Finite Difference Methods for Partial Differential Equations, by G. E. Forsythe and W. R. Wasow, 1960; 444 pages. (London: John Wiley & Sons Ltd.; New York: John Wiley & Sons). The Computer Journal, 4(1):78–79, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/4/1/78.full.pdf+

html. From the first paragraph: “This is the first serious book, in the era of automatic digital computing machines, which is devoted solely to partial differential equations and which treats all three major types. Its publication is therefore a major event, particularly at this state of intense development in the numerical analysis of partial differential equations.”.

Fox:1961:CMT

[Fox61b]  L. Fox.  Computing ma-

chines for teaching and research. The Computer Journal, 4(3):212–216, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/4/3/212.full.pdf+
html;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_03/040212.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/
212.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_03/tiff/
213.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_03/tiff/
214.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_03/tiff/
215.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_03/tiff/
216.tif.

Fox:1961:SCA

[Fox61c] L. Fox. Some comments on the accuracy and convenience of finite-difference processes in scientific computation. Proc. Ass. France de Calcul, ??(??):63–73, ???? 1961.

Fox:1962:CMO

[Fox62a] L. Fox. Chebyshev methods for ordinary differential equations. The Computer Journal, 4(4):318–331, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/4/4/318.full.pdf+

html;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/040318.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/
318.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
319.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
320.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
321.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
322.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
323.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
324.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
325.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
326.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
327.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
328.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
329.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/
330.tif;   http://www3.oup.
co.uk/computer_journal/hdb/

Volume_04/Issue_04/tiff/
331.tif.

Fox:1962:NSO

[Fox62b] L. Fox, editor. Numerical Solution of Ordinary and Partial Differential Equations. Based on a Summer School Held in Oxford, August–September 1961. Pergamon Press, New York, NY, USA, 1962. 509 pp. LCCN QA371 .L758 1962.

Fox:1963:PDE

[Fox63] L. Fox. Partial differential equations. The Computer Journal, 6(1):69–74, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/6/1/69.full.pdf+
html;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/060069.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/
69.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/
70.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/
71.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/
72.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/
73.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/
74.tif.

Fox:1964:INL

[Fox64] L. Fox. An Introduction to Numerical Linear Algebra. Monographs on Numerical Analysis. Clarendon Press, Oxford, UK, 1964. xi + 295 pp. LCCN QA251 .F65. Reprinted in 1967.

Fox:1965:CLS

[Fox65a] L. Fox. Chebyshev least squares approximation. In D. C. Handscomb, editor, Approximation [Computing Laboratory Summer School], chapter 5, page ?? Pergamon Press, New York, NY, USA, 1965.

Fox:1965:DPC

[Fox65b] L. Fox. Determination and properties of Chebyshev expansions. In D. C. Handscomb, editor, Approximation [Computing Laboratory Summer School], chapter 6, page ?? Pergamon Press, New York, NY, USA, 1965.

Fox:1965:INL

[Fox65c] L. Fox. An Introduction to Numerical Linear Algebra, with Exercises. Monographs on Numerical Analysis. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1965. xi + 327 pp. LCCN QA251 .F65 1965.

Fox:1965:LSA

[Fox65d] L. Fox. Least squares approximation, orthogonal polynomials. In D. C. Handscomb, editor, Approximation [Computing Laboratory Summer School], chapter 4, page ?? Pergamon Press, New York, NY, USA, 1965.

Fox:1965:NMA

[Fox65e] L. Fox. Numerical methods for arch dams. In ????, editor, Proceedings of the International Symposium, Southampton, pages 169–187. Pergamon Press, New York, NY, USA, 1965.

Fox:1965:PAP

[Fox65f]  L. Fox.  Presidential Ad-

dress: The proper use of recurrence relations. Mathematical Gazette, 49(370):371– 387, December 1965. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.
jstor.org/stable/3612171.

Fox:1966:APN

[Fox66a] L. Fox, editor. Advances in Programming and Non-Numerical Computation, volume 3 of Proceedings of summer schools organised by the Oxford University Computing Laboratory and the Delegacy for Extra-mural Studies. Pergamon Press, New York, NY, USA, 1966. viii + 218 pp. LCCN QA76 .F64 1966.

Fox:1966:GS

[Fox66b] L. Fox. General survey. In Walsh [Wal66], page ?? LCCN QA297 .N84 1967.

Fox:1967:RIC

[Fox67] L. Fox. Romberg integration for a class of singular integrands. The Computer Journal, 10(1):87–93, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/10/1/87.full.pdf+
html;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/100087.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/
87.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/
88.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/
89.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/
90.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/
91.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/
92.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_01/tiff/
93.tif.

Fox:1968:BRP

[Fox68a] L. Fox. Book review: Procédures Algol en Analyse Numérique, 324 pages. (Published by Min. de l’Éducation Nationale). The Computer Journal, 11(1):30, January 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/

content/11/1/30.full.pdf+
html.

Fox:1968:CSN

[Fox68b] L. Fox. Comments on singularities in numerical integration and the solution of differential equations. Colloquia Math Soc. Janos Bolyai, ??(??):61– 91, 1968. 3 - Numerical Methods, Tihany, Hungary.

Fox:1969:CEM

[Fox69a]  L. Fox.  Computer evalu-

ation of mathematical function. The Computer Journal, 12(3):267, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/12/3/267.full.pdf+
html.

Fox:1969:MPP

[Fox69b] L. Fox. Mathematical and physical polynomials. In Dejon and Henrici [DH69], pages 63–68. ISBN 0-471-20300-9. LCCN QA212 .C65.

Fox:1970:I

[Fox70] L. Fox. Introduction. In Hayes [Hay70], chapter 1, pages 1– 3. ISBN 0-485-11109-8. LCCN QA297 .N85 1970.

Fox:1971:HGM

[Fox71a] L. Fox. How to get meaningless answers in scientific computation (and what to do about it). IMA Bulletin, 7(10):296– 302, 1971.

Fox:1971:SES

[Fox71b] L. Fox. Some experiments with singularities in linear elliptic partial differential equations. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 323(1553):179–190, June 1971. CODEN PRLAAZ. ISSN 0962-8444 (print), 2053-9177 (electronic).

Fox:1972:TNA

[Fox72] L. Fox. The teaching of numerical analysis. Bulletin of the Institute of Mathematics and its Applications, 8(??):255–258, ???? 1972. CODEN IMTABW. ISSN 0950-5628.

Fox:1973:RNM

[Fox73a] B. L. Fox. Reducing the number of multiplikations in iterative processes. Acta Informatica, 3 (1):43–45, December 31, 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fox:1973:INL

[Fox73b] L. Fox. An Introduction to Numerical Linear Algebra. Monographs on Numerical Analysis. Clarendon Press, Oxford, UK, 1973. ISBN 0-19-853402-7 (hardcover), 0-19-853407-8 (paperback). xi + 328 pp. LCCN QA251 .F65. Reprinted in [Fox79b].

Fox:1974:BRI

[Fox74a] L. Fox. Book review: Introduction to Matrix Computations, G. W. Stewart, 1973; 441 pages (Academic Press). The Computer Journal, 17(3):200, March 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/17/3/200.full.pdf+
html.

Fox:1974:TFB

[Fox74b] L. Fox. Two free-boundary problems with singularities. Acta Universitatis Carolinae. Mathematica et Physica, 15(1– 2):19–23, 1974. ISSN 0001-7140.

Fox:1975:INA

[Fox75a] L. Fox. Introductory numerical analysis. In B. L. Hewitt et al., editors, Computational methods and problems in aeronautical fluid dynamics: Proceedings of IMA Conference, 1974, pages 15–52. Academic Press, New York, USA, 1975.

Fox:1975:WBN

[Fox75b] L. Fox. What are the best numerical methods? In J. R. Ockendon and W. R. Hodgkins, editors, Moving Boundary Problems in Heat flow and Diffusion, pages 210–241. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1975.

Fox:1976:BRC

[Fox76a] L. Fox. Book review: The Chebyshev Polynomials, Theodore J. Rivlin, 1974; 186 pages (John Wiley). The Computer Journal, 19(1):42, January 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        comjnl.oxfordjournals.org/
                                        content/19/1/42.full.pdf+
                                        html.

Fox:1976:NAC

[Fox76b] L. Fox. Numerical analysis, computers and problem solving. Interdisciplinary Science Reviews, 1(2):167–175, ???? 1976. CODEN ISCRD8. ISSN 0308-0188 (print), 1743-2790 (electronic).

Fox:1977:FDM

[Fox77a] L. Fox. Finite-difference methods for elliptic boundary-value problems. In Jacobs [Jac77], pages 799–881. ISBN 0-12-378650-9. LCCN QA297 .C646 1976.

Fox:1977:IEP

[Fox77b] L. Fox. An inverse eigenvalue problem. Bulletin of the Institute of Mathematics and its Applications, 13(6):162–163, ???? 1977. CODEN IMTABW. ISSN 0905-5628.

Fox:1978:AAJ

[Fox78] L. Fox. All about Jim Wilkinson, with a commemorative snippet on backward error analysis. In The Contribution of Dr. J. H. Wilkinson to Numerical Analysis, volume 19 of Symposium Proceedings Series, pages 1–20. The Institute of Mathematics and its Applications, Southend-On-Sea, Essex, UK, 1978.

Fox:1979:FDS

[Fox79a] L. Fox. Finite differences and singularities in elliptic problems. In I. Gladwell and R. Wait, editors, A Survey of Numerical Methods for Partial Differential Equations, chapter 3, pages 42–79. Clarendon Press, Oxford, UK, 1979.

Fox:1979:INL

[Fox79b] L. Fox. An Introduction to Numerical Linear Algebra. Monographs on Numerical Analysis. Clarendon Press, Oxford, UK, 1979. ISBN 0-19-853402-7, 0-19-853407-8. xi + 328 pp. LCCN QA251 .F65. Reprint of [Fox73b].

Fox:1979:SP

[Fox79c] L. Fox. The Stefan problem.

In I. Gladwell and R. Wait, editors, A Survey of Numerical Methods for Partial Differential Equations, chapter ??, pages 332–356. Clarendon Press, Oxford, UK, 1979.

Fox:1980:NMB

[Fox80]  L. Fox.  Numerical meth-

ods for boundary value problems. In I. Gladwell and D. K. Sayers, editors, Computational techniques for ordinary differential equations (Proceedings of the Conference University of Manchester, Manchester, 1978), Inst. Math. Appl. Conf. Ser., chapter 9, pages 175–216. Academic Press, New York, USA, 1980. ISBN 0-12-285780-1.

Fox:1981:FDM

[Fox81] L. Fox. Finite-difference methods for elliptic boundary value problems. Yingyong Shuxue yu Jisuan Shuxue, (3):11–34, 1981. Translated from the English by Xue Zong Liu and Lie Heng Wang.

Fox:1983:CSL

[Fox83] L. Fox. Computers still have a lot to learn about good numerical analysis and numerical algorithms. NAG Newsletter, 2(??): 5–15, ???? 1983. ISSN 0269-0780.

Fox:1984:NAH

[Fox84] L. Fox. Numerical analysis in higher education. Bulletin of the Institute of Mathematics and its Applications, 20(??): 103–108, ???? 1984. CODEN IMTABW. ISSN 0950-5628. Presented at the Symposium on “The contribution of Leslie Fox to numerical analysis”.

Fox:1985:BNA

[Fox85a] L. Fox. The beginnings of numerical analysis. Lecture to British Society for the History of Mathematics, King’s College, London, 1985.

Fox:1985:CAM

[Fox85b] L. Fox. Chairman’s address: Mathematical and numerical modelling. In ????, editor, Proceedings of the Tenth Canadian Congress of Applied Mechanics, University of Western Ontario, pages 1–10. ????, ????, 1985.

Fox:1987:ENA

[Fox87a] L. Fox. Early numerical analysis in the United Kingdom. In Crane [Cra87], pages 21– 39. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

Fox:1987:JHW

[Fox87b] L. Fox. James Hardy Wilkinson, 27 September 1919–5 October 1986. Biographical Memoirs of Fellows of the Royal Society, 33:669–708, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

URL http://www.jstor.org/
stable/7699a67;   https:/
/royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.
1987.0024.

Fox:1987:ODJ

[Fox87c] Leslie Fox. Obituary: Dr.

J. H. Wilkinson, F.R.S. The Computer Journal, 30(1): 1, 1987. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/1.
tif.

Fox:1990:ENA

[Fox90a] L. Fox. Early numerical analysis in the United Kingdom. In Nash [Nas90], pages 280–300. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Fox:1990:EJW

[Fox90b] L. Fox. Epilogue. Jim Wilkinson: Some after-dinner sentiments. In Cox and Hammarling [CH90], pages 329–335. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$45.00.

Fox:1990:NST

[Fox90c] L. Fox. The Numerical Solution of Two-point Boundary Problems in Ordinary Differential Equations. Monographs on Numerical Analysis. Dover Publications, Inc., New York, NY, USA, 1990. ISBN 0-486-66495-3. xii + 371 pp. Reprint of [Fox57a].

Fox:1993:NA

[Fox93] L. Fox. Numerical analysis. In P. G. Drazin, editor, Collected papers of Lewis Fry Richardson, page ?? Cambridge University Press, Cambridge, UK, 1993. This is Leslie Fox’s final publication.

Fox:1968:CPN

[FP68] L. Fox and I. B. Parker. Chebyshev Polynomials in Numerical Analysis. Oxford mathematical handbooks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1968. ISBN 0-19-859614-6. ix + 205 pp. LCCN QA297 .F65. Reprinted in 1972 with corrections, but same ISBN.

Fox:1962:MCA

[FR62] L. Fox and H. F. Rainsford. Mechanical computing. Appendix I. In D. Clark, editor, Plane and geodetic surveying, pages 615– 634. Constable, ????, 1962.

Fox:1941:NUR

[FS41] L. Fox and R. V. Southwell.

Note on the use of relaxation methods for the computation of two-dimensional stress systems. Report E. F. 466, Aeronautical Research Committee, Structure subcommittee, Strut 512, ????, 1941. This is Leslie Fox’s first publication.

Fox:1942:STB

[FS42] L. Fox and R. V. Southwell.

On the stressing of tapered box beams. Report ??, Aeronautical Research Committee, Structure subcommittee 6034, Strut 598, ????, 1942.

Fox:1944:FGT

[FS44] L. Fox and R. V. Southwell.

On the flow of gas through a nozzle with velocities exceeding the speed of sound. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 183 (992):38–54, August 1944. CO-DEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Fox:1945:RMA

[FS45] L. Fox and R. V. Southwell.

Relaxation methods applied to engineering problems. VII A. Biharmonic analysis as applied to the flexure and extension of flat elastic plates. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 239 (810):419–460, October 1945. CODEN PTRMAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Fox:1946:SHT

[FS46] L. Fox and R. V. Southwell. On the stresses in hooks, and their determination by relaxation methods. Inst Mech Eng Appl Mech Proc, 155(??):1–19, ???? 1946.

Fox:1969:BSL

[FS69] L. Fox and R. Sankar. Boundary singularities in linear elliptic differential equations. Journal of the Institute of Mathematics and its Applications, 5(3):340–350, September 1969. CODEN JMTAA8. ISSN 0020-2932. URL https:

                                        //academic.oup.com/imamat/
                                        article/5/3/340/697349.

Fox:1973:RFM

[FS73] L. Fox and R. Sankar. The Regula-Falsi method for freeboundary problems. Journal of the Institute of Mathematics and its Applications, 12(1):49– 54, August 1973. CODEN JM-TAA8. ISSN 0020-2932.

Fox:1980:SEI

[FV80] L. Fox and M. R. Valenca. Some experiments with interval methods for two-point boundary value problems in ordinary differential equations. BIT (Nordisk tidskrift for informationsbehandling), 20(1):67–82, March 1980. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0006-
3835&volume=20&issue=1&spage=
67.

Fox:1966:FCU

[FW66] L. Fox and J. E. Walsh. Further comments on University work. In Walsh [Wal66], page ?? LCCN QA297 .N84 1967.

Fox:1975:F

[FW75] L. Fox and J. H. Wilkinson.

Foreward. In The NAG Library Manual, pages 1–3. The Numerical Algorithms Group Ltd, Wilkinson House, Jordan Hill Road, Oxford OX2 8DR, UK, 1975.

Fox:1963:C

[FY63] L. Fox and Andrew Young.

Correspondence. The Computer Journal, 5(4):296, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/
content/5/4/296.full.pdf+
html.

Gauss:1958:FVL

[Gau58]  F. G. Gauss.  Fünfstellige

vollständige logarithmische und trigonometrische Tafeln. Herausgegeben von Dr.-Ing. H. H. Gobbin. (German) [Fivedigit complete logarithmic and trigonometric tables. Published by Dr.-Ing. H. H. Gobbin]. Verlag Konrad Wittwer, Stuttgart, West Germany, 1958. xxxii + 184 pp. 391.-400. Aufl.

Gautschi:1975:OCV

[Gau75] Walter Gautschi. Optimally conditioned Vandermonde matrices. Numerische Mathematik, 24(1):1–12, February 1975. CO-DEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Gear:1965:HMI

[Gea65] C. W. Gear. Hybrid methods for initial value problems in ordinary differential equations. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(1):69–86, 1965. ISSN 0887-459X (print), 2168-3581 (electronic).

Green:1965:SSC

[GM65] D. J. Green and S. Michaelson. Series solution of certain Sturm–Liouville eigenvalue problems. The Computer Journal, 7(4):322–336, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Guderley:1966:RTP

[GN66] Karl G. Guderley and Paul J. Nikolai. Reduction of twopoint boundary value problems in a vector space to inital value problems by projection. Numerische Mathematik, 8(3):

270–289, May 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goldfarb:1972:MMI

[Gol72]  D. Goldfarb.  Modification

methods for inverting matrices and solving systems of linear algebraic equations. Mathematics of Computation, 26(120): 829–852, October 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Goodwin:1965:NPC

[Goo65] E. T. Goodwin. Note on a paper by S. C. R. Dennis. Mathematical Proceedings of the Cambridge Philosophical Society, 61 (1):299, January 1965. CO-DEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Glowinski:1980:EPS

[GPP80] R. Glowinski, J. Periaux, and O. Pironneau. An efficient preconditioning scheme for iterative numerical solutions of partial differential equations. Applied Mathematical Modelling, 4 (3):187–192, June 1980. CO-DEN AMMODL. ISSN 0307-904x (print), 1872-8480 (electronic).

Grad:1971:MB

[Gra71] J. Grad. Matrix balancing. The Computer Journal, 14(3):280– 284, March 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Gunzenhauser:1967:BRA

[Gun67] R. Gunzenhäuser. Book report: Advances in Programming and Non-Numerical Computations, Herausgegeben von L. Fox. viii, 218 S. Oxford-London-Edinburgh-New York-Paris-Frankfurt: Pergamon Press. 1966. Computing, 2 (2):180, June 1967. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Griffin:1963:NSP

[GV63] D. S. Griffin and R. S. Varga. Numerical solution of plane elasticity problems. Journal of the Society for Industrial and Applied Mathematics, 11 (4):1046–1062, December 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (elec-

tronic). URL http://www.
math.kent.edu/~varga/pub/
paper_24.pdf. See remark

[Bue66].

Griffin:1966:CBR

[GV66] D. S. Griffin and R. S. Varga. Comments on Bueckner’s Remark on a Paper by D. S. Griffin and R. S. Varga. SIAM Journal on Applied Mathematics, 14 (2):415–416, March 1966. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [GV63, Bue66].

Gyr:1974:EPM

[Gyr74] Marcel Gyr. Eigenvalues of partitioned matrices “generated” by a matrix with known eigenvalues. IEEE Transactions on Automatic Control, AC-19:88– 89, 1974. CODEN IETAA9. ISSN 0018-9286 (print), 1558-2523 (electronic).

Haselgrove:1961:SNL

[Has61] C. B. Haselgrove. The solution of non-linear equations and of differential equations with two-point boundary conditions. The Computer Journal, 4(3): 255–259, March 1961. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Hayes:1970:NAF

[Hay70] J. G. Hayes, editor. Numerical Approximation to Functions and Data. Athlone Press, London, UK, 1970. ISBN 0-485-11109-8. viii + 177 pp. LCCN QA297 .N85 1970.

Hegedus:1975:LEI

[Heg75] Cs. J. Hegedűs. Laurent expansion of an inverse of a function matrix. Periodica Mathematica Hungarica, 6(1):75–86, March 1975. CODEN PMH-GAW. ISSN 0031-5303 (print), 1588-2829 (electronic).

Henrici:1959:BRN

[Hen59]  Peter Henrici.  Book re-

view: The Numerical Solution of Two-Point Boundary Problems in Ordinary Differential Equations, by L. Fox. Mathematical Tables and Other Aids to Computation, 13(68): 322–323, October 1959. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

                                        tronic). URL https://www.
                                        jstor.org/stable/2002815.

Hersch:1958:ROS

[Her58] Joseph Hersch. Récurrences d’ordre supérieur pour des équations aux différences. (French) [Higher-order repetitions for difference equations]. Comptes rendus de l’Académie des sciences, Paris, 246:264–367, 1958. ISSN 0001-4036 (print), 2419-6304 (electronic).

Hildebrand:1956:INA

[Hil56] F. B. Hildebrand. Introduction to Numerical Analysis. Mc-Graw-Hill, New York, NY, USA, 1956. x + 511 pp.

Hitotumatu:1974:NMC

[Hit74]  Sin Hitotumatu.  A new

method for the computation of square root, exponential and logarithmic functions through hyperbolic CORDIC. Revue d’Analyse Numérique et de la Théorie de l’Approximation, 3(2):173– 180, 1974. ISSN 1010-3376 (print), 2457-8118 (electronic).

                                        URL https://ictp.acad.ro/
                                        jnaat/journal/article/view/
                                        1974-vol3-no2-art7.

Hof:1959:TPN

[Hof59] Hans Hof. Ten Place Natural Trigonometric Tables: Sine– Tangent, 0 to 90 Degrees. Professional Supply Company, Jenkintown, PA, USA, 1959. iv + 1084 pp.

Holnapy:1974:NAD

[Hol74] D. Holnapy. Numerical algorithm to determine the difference operators of boundaryvalue problems. Acta Tech. Acad. Sci. Hungar., 76:413–423, 1974. ISSN 0001-7035.

Horgan:1956:TCP

[Hor56] R. B. Horgan. Table of coefficients for the partial summation of series. Mathematical Tables and Other Aids to Computation, 10(55):156–162, July 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Householder:1955:GED

[Hou55] Alston S. Householder. The generation of error in digital computation. Technical Report ORNL-1983, Oak Ridge National Laboratory, Knoxville, TN, USA, October 11, 1955. iv + 79 pp.

URL https://www.osti.gov/
biblio/4380237-generation-
error-digital-computation.

Hwang:1973:RRS

[HT73] W. G. Hwang and John Todd.

A recurrence relation for the square root. Journal of Approximation Theory, 9(4):299–306, December 1973. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Huckel:1956:TFW

[Huc56] Vera Huckel. Tabulation of the f_{λ} functions which occur in the aerodynamic theory of oscillating wings in supersonic flow. NACA Technical Note, 1956 (3606):59, 1956. ISSN 0097-2177.

Hammer:1960:QFI

[HW60] Preston C. Hammer and Howard H. Wicke. Quadrature formulas involving derivatives of the integrand. Mathematics of Computation, 14(69): 3–7, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://links.jstor.
org/sici?sici=0025-5718(
196001)14:69<3:QFIDOT>2.0.
CO;2-O&origin=MSN.

Hsiao:1977:GMC

[HW77] George C. Hsiao and Wolfgang Wendland. On Galerkin’s method for a class of integral equations of the first kind. Applicable Analysis, 6(2):155– 157, January 1977. CODEN APANCC. ISSN 0003-6811.

Ilioi:1966:EER

[Ili66] C. Ilioi. Sur l’évaluation de l’erreur dans la résolution des systèmes linéaires. Analele Ştiinţifice ale Universităţii “Al. I. Cuza” din Iasi. Secţiunea I a. Matematică. Serie Nouă, 12: 313–320, 1966. ISSN 0041-9109.

Iri:1968:MMD

[Iri68]  Masao Iri.  A method of

multi-dimensional linear interpolation. Information Processing in Japan, 8:33–36, 1968. ISSN 0537-6246.

Iwata:1970:AFM

[Iwa70]  Giiti Iwata.  Approximate

formulas for the mean value of a function over a tetrahedron. Natural Science Report of the Ochanomizu University, 21:39–47, 1970. CODEN NASOA5. ISSN 0029-8190.

URL https://teapot.lib.
ocha.ac.jp/records/34934.

Jacobs:1977:SAN

[Jac77] David A. H. Jacobs, editor.

The State of the Art in Numerical Analysis: Proceedings of the Conference on the State of the Art in Numerical Analysis held at the University of York, April 12th–15th, 1976. Academic Press, New York, USA, 1977. ISBN 0-12-378650-9. LCCN QA297 .C646 1976.

Jaeger:1956:NVT

[Jae56] J. C. Jaeger. Numerical values for the temperature in radial heat flow. Journal of Mathematical Physics, 34:316–321, 1956. CODEN JMAPAQ. ISSN 0097-1421.

Jeltsch:1976:NSM

[Jel76] Rolf Jeltsch. Note on Astability of multistep multiderivative methods. BIT (Nordisk tidskrift for informationsbehandling), 16(1):74–78, March 1976. CODEN BITTEL, NBITAB. ISSN 0901-246X.

Jenson:1959:VFR

[Jen59] V. G. Jenson. Viscous flow round a sphere at low Reynolds numbers (< 40). Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 249 (1258):346–366, January 1959. CODEN PRLAAZ. ISSN 0962-8444 (print), 2053-9177 (electronic).

Jones:1961:NSC

[Jon61] J. G. Jones. On the numerical solution of convolution integral equations and systems of such equations. Mathematics of Computation, 15(74):131–142, April 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kaye:1955:TFE

[Kay55] Joseph Kaye. A table of the first eleven repeated integrals of the error function. Journal of Mathematical Physics, 34:119– 125, 1955. CODEN JMAPAQ. ISSN 0097-1421.

Keeping:1971:CAT

[Kee71] A. J. Keeping. On certain almost-triangular band matrices. SIAM Journal on Numerical Analysis, 8(2):288–303, June 1971. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kennebrock:1976:DSN

[Ken76] W. Kennebrock. A difference scheme for the numerical solution of the Boltzmannequation. In Methoden und Verfahren der mathematischen Physik, Band 16, pages 91– 103. Bibliographisches Institut, Mannheim, Germany, 1976.

Kershaw:1987:AMF

[Ker87] D. Kershaw. An analysis of the method of L. Fox and L. Hayes for the factorization of a polynomial. Linear Algebra and its Applications, 86(??):179–187, February 1987. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
002437958790293X. See [FH68].

Klerer:1968:ERT

[KG68] Melvin Klerer and Fred Grossman. Error rates in tables of indefinite integrals. Industrial Mathematics, 18:31–62, 1968. CODEN IMTHAI. ISSN 0019-8528.

Kagiwada:1978:IMI

[KK78] H. Kagiwada and R. Kalaba. Imbedding methods for integral equations with applications. Journal of Optimization Theory and Applications, 24(1):29–57, 1978. CO-DEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic). Issue devoted to numerical methods for solving integral equations.

Krarup:1956:MMM

[KS56] T. Krarup and Bj. Svejgaard.

A method for matrix multiplication, matrix inversion, and problems of adjustment by punched card equipment. Geodætisk Institut, København. Meddelelser, 1956(31):31, 1956.

Kuki:1972:AAC

[Kuk72a] Hirondo Kuki. ACM Algorithm 421: Complex gamma function with error control [S14]. Communications of the ACM, 15 (4):271–272, April 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuki:1972:CGF

[Kuk72b] Hirondo Kuki. Complex gamma function with error control [S14]. Communications of the ACM, 15(4):262–267, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kulikov:1973:EEA

[Kul73] È. L. Kulikov. Estimating errors in the approximate solution of boundary value problems in electrodynamics by the variational method. Radio Engineering and Electronic Physics. A translation of Radiotekhnika i Èlektronika, a U.S.S.R. Academy of Sciences Publication, 18:926–929, 1973. ISSN 0033-7889.

Kuntzmann:1958:CFL

[Kun58]  J. Kuntzmann.  Calcul de

fonctions de Legendre. (French) [Calculation of Legendre functions]. In Symposium on modern computers: Proceedings of the Rome Symposium (17– 18 October 1956) organized by the Preparatory Committee of the International Computation Center under the auspices of UNESCO, pages 1–3. Veschi, Rome, Italy, 1958.

Kundu:1977:AIN

[Kun77]  M. C. Kundu.  Approxi-

mate integration near a simple pole using Chebyshev points of the second kind. Bulletin Mathématique de la Société des Sciences Mathématiques de la République Socialiste de Roumaine. Nouvelle Série, 21(69)(3–4):323–330, 1977. ISSN 0007-4691. URL https://www.

jstor.org/stable/43759449.

Kurtz:1976:SFO

[KWC76] Lawrence A. Kurtz, David G.

Weinman, and Alan K. Cline. The solution of a fourth-order partial differential equation using the method of lines. Applicable Analysis, 5(3):201– 206, January 1976. CODEN APANCC. ISSN 0003-6811.

L:1961:BRF

[L.61] Y. L. L. Book review: L.

Fox, Tables of Weber Parabolic Cylinder Functions and Other Functions for Large Arguments, National Physical Laboratory Mathematical Tables Volume 4, Her Majesty’s Stationery Office, London, 1960, iii + 40 p., 28 cm. (Paperback). Mathematics of Computation, 15(75): 310–311, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                        URL http://www.jstor.org/
                                        stable/2002928;   https://
                                        www.ams.org/journals/mcom/
                                        1961-15-075/S0025-5718-61-
                                        99215-8/S0025-5718-61-99215-
                                        8.pdf.

Lado:1971:NFT

[Lad71] F. Lado. Numerical Fourier transforms in one, two, and three dimensions for liquid state calculations. Journal of Computational Physics, 8:417–433, 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Langer:1959:NAP

[Lan59] R. E. Langer, editor. On Numerical Approximation. Proceedings of a Symposium, Madison, April 21–23, 1958. The University of Wisconsin Press, Madison, WI, USA, 1959. x + 462 pp. LCCN QA3 .U45 no. 1. Publication no. 1 of the Mathematics Research Center, U.S. Army, the University of Wisconsin.

Ledsham:1961:NIU

[Led61] F. C. Ledsham. Numerical integration using sums of exponential functions. Mathematics of Computation, 15(73):48–51, January 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lemeire:1976:EMO

[Lem76]  Frans Lemeire.  Equilibra-

tion of matrices to optimize backward numerical stability. BIT (Nordisk tidskrift for informationsbehandling), 16(2):143– 145, June 1976. CODEN BIT-TEL, NBITAB. ISSN 0901-246X.

Lieberstein:1968:CNA

[Lie68]  H. Melvin Lieberstein.  A

Course in Numerical Analysis. Harper & Row, New York, NY, USA, 1968. xiv + 258 pp.

Lister:1960:NSH

[Lis60] Mary Lister. The numerical solution of hyperbolic partial differential equations by the method of characteristics. In Mathematical methods for digital computers, pages 165–179. Wiley, New York, NY, USA, 1960.

Longsine:1980:SRQ

[LM80] D. E. Longsine and S. F.

McCormick. Simultaneous Rayleigh-quotient minimization methods for Ax = λBx. Linear Algebra and its Applications, 34(??):195–234, December 1980. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0024379580901664.

Lotkin:1956:NSL

[Lot56] Mark Lotkin. Note on the sensitivity of least squares solutions. Journal of Mathematical Physics, 35:309–311, 1956. CODEN JMAPAQ. ISSN 0097-1421.

Lipow:1956:RE

[LZ56] M. Lipow and S. A. Zwick. On the roots of the equation: Y_{1}(mx)[xJ_{1}(x) − BJ_{0}(x)] − J_{1}(mx)[xY_{1}(x) − BY_{0}(x)] = 0.. Journal of Mathematical Physics, 34:308–315, 1956. CO-DEN JMAPAQ. ISSN 0097-1421.

Mayers:1964:DAL

[May64] D. F. Mayers. The deferred approach to the limit in ordinary differential equations. The Computer Journal, 7(1):54–57, January 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

McClamroch:1985:BRG

[McC85] N. Harris McClamroch. Book review: A Guide to Simulation. By Paul Bratley, Bennett L. Fox, and Linus E. Schrage. Springer-Verlag, New York, 1.983. xviv + 383 pp. ISBN 0-387-90820-X. SIAM Review, 27(2):292–293, June 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McGee:1967:BRA

[McG67] William C. McGee. Book review: Advances in Programming & Non-Numerical Computation Edited by L. Fox. Pergamon Press, Inc., New York, 1966. viii + 218 pp. SIAM Review, 9(1):133–136, January 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McKee:2018:MRP

[McK18] Sean McKee. My recollections of Professor Leslie Fox. Mathematics Today, 54(4):148, 2018. ISSN 1361-2042.

Mendelsohn:1956:SEP

[Men56]  N. S. Mendelsohn.  Some

elementary properties of ill conditioned matrices and linear equations. American Mathematical Monthly, 63(5): 285–295, May 1956. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.
jstor.org/stable/2310510.

Meyer:1957:MIL

[MH57] H. I. Meyer and B. J.

Hollingsworth. A method of inverting large matrices of special form. Mathematical Tables and Other Aids to Computation, 11 (58):94–97, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Morris:1977:CPN

[MH77] A. G. Morris and T. S. Horner.

Chebyshev polynomials in the numerical solution of differential equations. Mathematics of Computation, 31(140):881–891, October 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miller:1949:CMT

[Mil49] J. C. P. Miller. The construction of mathematical tables. The scientific journal of the Royal College of Science, 20 (??):1–11, ???? 1949. CODEN SJRCAE. ISSN 0371-0467.

Milne:1953:NSD

[Mil53] William Edmund Milne. Numerical Solution of Differential Equations. Wiley, New York, NY, USA, 1953. xii + 275 pp.

Miller:1959:EAS

[Mil59] J. C. P. Miller. Extremal approximations — a summary. In On numerical approximation. Proceedings of a Symposium, Madison, April 21-23, 1958, volume no. 1 of Publication of the Mathematics Research Center, U.S. Army, the University of Wisconsin, pages 329– 340. The University of Wisconsin Press, Madison, WI, USA, 1959. Edited by R. E. Langer.

Miller:1960:TGE

[Mil60] G. F. Miller. Tables of Generalized Exponential Integrals. Her Majesty’s Stationery Office, London, UK, 1960. iii + 43 pp.

Munz:1981:UEC

[Mun81] Harry Munz. Uniform expansions for a class of finite difference schemes for elliptic boundary value problems. Mathematics of Computation, 36(153): 155–170, January 1981. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mayers:1999:OLF

[MW99] David F. Mayers and J. E.

Walsh. Obituary: Leslie Fox. Bulletin of the London Mathematical Society, 31(2):241–247, March 1999. CODEN LMS-BBT. ISSN 0024-6093 (print), 1469-2120 (electronic).

Nash:1990:HSC

[Nas90] Stephen G. Nash, editor. A History of Scientific Computing. ACM Press history series. Addison-Wesley and ACM Press, Addison-Wesley and New York, NY 10036, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990.

Nedyalkov:1970:ATI

[Ned70]  I. P. Nedyalkov.  An ap-

proach in the theory of incorrect problems. Comptes rendus de l’Académie bulgare des sciences, 23:5–8, 1970. CODEN DBANAD. ISSN 0366-8681.

Nichols:1969:GCO

[NF69] Nancy K. Nichols and L. Fox.

Generalized consistent ordering and the optimum successive over-relaxation factor. Numerische Mathematik, 13(5): 425–433, October 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nicolovius:1960:ISB

[Nic60]  Rüdiger Nicolovius.  Iter-

ative solution of boundary value problems using remainder terms of Taylor expansions. In Symposium on the numerical treatment of ordinary differential equations, integral and integro-differential equations (Rome, 1960), pages 162– 173. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1960.

Nickel:1967:ERP

[Nic67] Karl L. E. Nickel. Extension of a recent paper by Fox, Henrici and Moler on eigenvalues of elliptic operators. SIAM Journal on Numerical Analysis, 4(4): 483–488, December 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nichols:1993:OPL

[Nic93] Nancy K. Nichols. Obituary: Professor Leslie Fox. Bulletin of the Institute of Mathematics and its Applications, 29(11– 12):175–176, 1993. CODEN IMTABW. ISSN 0905-5628.

Nakashima:1976:NMN

[NOT76] Katsuya Nakashima, Katsushi Ohmori, and Kunio Tsuruta. The Newton method for nonlinear Volterra integro-differential equations. Bulletin of Science and Engineering Research Laboratory. Waseda University, (73):54–63, 1976. ISSN 0372-7181.

Olaofe:1977:TMS

[Ola77] G. Oluremi Olaofe. On the Tchebyschev method of solution of ordinary differential equations. Journal of Mathematical Analysis and Applications, 60(1):1–7, August 1977. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Orlov:1974:FTN

[Orl74a] K. Orlov. Forming of the table of numerical values of the polynomial by two operations for each new value. Mathematica Balkanica, 4:473–476, 1974. ISSN 0350-2007.

Orlov:1974:NKN

[Orl74b] K. Orlov. A new kind of numerical tables of functions made by mathematical spectra. Mat. Vesnik, 11/26:191–201, 1974. ISSN 0025-5165 (print), 2406-0682 (electronic). URL https:

//gdz.sub.uni-goettingen.
de/id/PPN311571026_0026.

Ortiz:1974:CPL

[Ort74] Eduardo L. Ortiz. Canonical polynomials in the Lanczos tau method. In Studies in numerical analysis (papers in honour of Cornelius Lanczos on the occasion of his 80th birthday), pages 73–93. Royal Irish Academy, Dublin, Ireland, 1974.

Ortiz:1981:OAT

[OS81] E. L. Ortiz and H. Samara.

An operational approach to the tau method for the numerical solution of nonlinear differential equations. Computing, 27 (1):15–25, March 1981. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Osborne:1960:EEP

[Osb60] M. R. Osborne. h^{2}-extrapolation in eigenvalue problems. Quarterly Journal of Mechanics and Applied Mathematics, 13(2): 156–168, 1960. CODEN QJM-MAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Osborne:1964:MFD

[Osb64a] M. R. Osborne. A method for finite-difference approximation to ordinary differential equations. The Computer Journal, 7(1):58–65, January 1964. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Osborne:1964:NMS

[Osb64b] M. R. Osborne. A new method for the solution of eigenvalue problems. The Computer Journal, 7(3):228–232, March 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Ockendon:2018:LFO

[OW18] John Ockendon and Andy Wathen. Leslie Fox and Oxford. Mathematics Today, 54(4):146– 147, 2018. ISSN 1361-2042.

P:1964:BRF

[P.64] H. P. Book review: L. Fox, Numerical Solution of Ordinary and Partial Differential Equations, Addison-Wesley Publishing Company, Inc., Reading, Mass., 1962, ix + 509 p., 23.5 cm. Mathematics of Computation, 18(87):528, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

jstor.org/stable/2003801.

Paige:1980:AEL

[Pai80] C. C. Paige. Accuracy and effectiveness of the Lanczos algorithm for the symmetric eigenproblem. Linear Algebra and its Applications, 34(??):235–258, December 1980. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0024379580901676.

Pace:1974:EALa

[PB74a] I. S. Pace and S. Barnett.

Efficient algorithms for linear system calculations. I. Smith form and common divisor of polynomial matrices. International Journal of Systems Science, 5:403–411, May 1974. CO-DEN IJSYA9. ISSN 0020-7721 (print), 1464-5319 (electronic).

Pace:1974:EALb

[PB74b] I. S. Pace and S. Barnett.

Efficient algorithms for linear system calculations. II. Minimal realization. International Journal of Systems Science, 5 (5):413–424, May 1974. CO-DEN IJSYA9. ISSN 0020-7721 (print), 1464-5319 (electronic).

Pereyra:1965:CDM

[Per65] Victor Pereyra. The correction difference method for nonlinear boundary value problems of class M. Revista de la Unión Matemática Argentina, 22(4): 184–201, 1965. ISSN 0041-6932 (print), 1669-9637 (electronic).

Pereyra:1967:ACD

[Per67] Victor Pereyra. Accelerating the convergence of discretization algorithms. SIAM Journal on Numerical Analysis, 4(4): 508–533, December 1967. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pelios:1972:MEC

[PK72] A. Pelios and R. W. Klopfenstein. Minimal error constant numerical differentiation (N.D.) formulas. Mathematics of Computation, 26(118): 467–475, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pollard:1964:RPPa

[Pol64] H. Pollard. Recent publications and presentations: Numerical Solution of Ordinary and Partial Differential Equations, by L. Fox. American Mathematical Monthly, 71 (8):938, October 1964. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL https://www.
                                        jstor.org/stable/2312444.

Rabinowitz:1960:AWL

[Rab60] Philip Rabinowitz. Abscissas and weights for Lobatto quadrature of high order. Mathematics of Computation, 14(69):47–52, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

links.jstor.org/sici?sici=
0025-5718(196001)14:69<47:
AAWFLQ>2.0.CO;2-Z&origin=
MSN.

Rainey:1969:CST

[Rai69] John W. Rainey. On comparatively stable tridiagonalization methods. Numerische Mathematik, 13(4):316–322, August 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rall:1978:RKG

[Ral78] L. B. Rall. Resolvent kernels of Green’s function kernels and other finite-rank modifications of Fredholm and Volterra kernels. Journal of Optimization Theory and Applications, 24(1):59–88, 1978. CO-DEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic). Issue devoted to numerical methods for solving integral equations.

Reynolds:1954:TSC

[Rey54] George E. Reynolds. Table of Squares of Cosecants. Air Force Cambridge Research Center, Electronics Research Directorate, Cambridge, MA, USA, 1954. iv + 89 pp.

Romanowski:1970:IMN

[RI70] M. Romanowski and L. Isaacs. Integration of modulated normal functions. Bulletin Géodésique (N.S.), (98):299– 308, 1970. ISSN 0007-4632.

Richardson:1911:IAA

[Ric11] Lewis Fry Richardson. IX. The approximate arithmetical solution by finite differences of physical problems involving differential equations with an application to the stresses in a masonry dam. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, A210(459–470):307– 357, January 1, 1911. CO-DEN PTRMAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic). URL https:

//royalsocietypublishing.
org/doi/10.1098/rsta.1911.
0009.

Reuss:1961:SRR

[RIY61] E. Reuss, R. Irving, and C. S. Yang. Some reflections on the relaxation of biharmonic differential equations in polar coordinates. Periodica Polytechnica. Engineering. Periodica Polytechnica. Maschinen- und Bauwesen, 5:335–344, 1961. ISSN 0031-5338.

Rheinboldt:1977:EMP

[RMF77] W. C. Rheinboldt, C. K. Mesztenyi, and J. M. Fitzgerald. On the evaluation of multivariate polynomials and their derivatives. BIT (Nordisk tidskrift for informationsbehandling), 17(4):437–457, December 1977. CODEN BITTEL, NBITAB. ISSN 0901-246X.

Robertson:1956:PCN

[Rob56] H. H. Robertson. Phase calculations for nuclear scattering on the Pilot ACE. Mathematical Proceedings of the Cambridge Philosophical Society, 52 (3):538–545, June 1956. CO-DEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Rossiter:1958:ARM

[Ros58] J. R. Rossiter. On the application of relaxation methods to oceanic tides. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 248(1255):482– 498, December 1958. CO-DEN PRLAAZ. ISSN 0962-8444 (print), 2053-9177 (electronic).

Rowe:1955:DAP

[Row55] Paul P. Rowe. Difference approximations to partial derivatives for uneven spacings in the network. Transactions. American Geophysical Union, 36(6): 995–1008, December 1955. CO-DEN TAGUAT. ISSN 0002-8606 (print), 2379-6723 (electronic).

Roth:1956:VMS

[RS56] J. P. Roth and D. S. Scott. A vector method for solving linear equations and inverting matrices. Journal of Mathematical Physics, 35:312–317, 1956. CODEN JMAPAQ. ISSN 0097-1421.

Rutishauser:1955:BNI

[Rut55] Heinz Rutishauser. Bemerkungen zur numerischen Integration gewöhnlicher Differentialgleichungen n-ter Ordnung. (German) [Remarks on the numerical integration of ordinary differential equations n-th order]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 6(6):497– 498, November 1955. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Stegun:1957:GBF

[SA57] Irene A. Stegun and Milton Abramowitz. Generation of Bessel functions on high speed computers. Mathematical Tables and Other Aids to Computation, 11(60):255–257, October 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1955:CZE

[Sal55] Herbert E. Salzer. Complex zeros of the error function. Journal of The Franklin Institute, 260(3):209–211, September 1955. CODEN JFINAB.

ISSN 0016-0032 (print), 1879-2693 (electronic).

Salzer:1956:FIO

[Sal56a] Herbert E. Salzer. Formulas for inverse osculatory interpolation. Journal of Research of the National Bureau of Standards (1934), 56:51–54, 1956. ISSN 0160-1741 (print), 2376-5259 (electronic).

Salzer:1956:FPS

[Sal56b] Herbert E. Salzer. Formulas for the partial summation of series. Mathematical Tables and Other Aids to Computation, 10 (55):149–156, July 1956. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1957:NIU

[Sal57] Herbert E. Salzer. Numerical integration of y′′ = ϕ(x, y, y′) using osculatory interpolation. Journal of The Franklin Institute, 263(5):401–409, May 1957. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

Salzer:1959:TOI

[Sal59] H. E. Salzer. Tables of Osculatory Interpolation Coefficients, volume 56 of National Bureau of Standards Applied Mathematics Series. United States Government Printing Office, Washington, DC, USA, 1959. xi + 25 pp.

Schiop:1978:APR

[Şch78a] A. I. Şchiop. An alternative proof of a result of Lees. Universitǎţii din Craiova. Analele. Matematică, Fizică-Chimie, 6: 43–45, 1978.

Schonfelder:1978:CEE

[Sch78b] J. L. Schonfelder. Chebyshev expansions for the error and related functions. Mathematics of Computation, 32(144):1232– 1240, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Scraton:1973:PAS

[Scr73] R. E. Scraton. Polynomial approximations to the solution of the heat equation. Mathematical Proceedings of the Cambridge Philosophical Society, 73 (1):157–165, January 1973. CO-DEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Selfridge:1958:RPN

[Sel58] R. G. Selfridge. Recent publications: The Numerical Solution of Two-Point Boundary Problems in Ordinary Differential Equations, by L. Fox. American Mathematical Monthly, 65 (8):641–642, October 1958. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL https://www.
                                        jstor.org/stable/2309147.

Shimizu:1973:NID

[Shi73] Tatsujiro Shimizu. Numerical integration of differential equations on 0 ≤ x < ∞. TRU Mathematics, 9:63–67, 1973. ISSN 0496-6597.

Shintani:1981:PRK

[Shi81] Hisayoshi Shintani. On pseudo-Runge–Kutta methods of the third kind. Hiroshima Mathematical Journal, 11(2):247–254, July 1981. ISSN 0018-2079.

URL http://projecteuclid.
org/euclid.hmj/1206134099.

Singer:1964:ENA

[Sin64] James Singer. Elements of Numerical Analysis. Academic Press Textbooks in Mathematics. Academic Press, New York, USA, 1964. x + 395 pp.

Salzer:1958:TBO

[SK58] Herbert E. Salzer and Genevieve M.

Kimbro. Tables for Bivariate Osculatory Interpolation over a Cartesian Grid. Convair Division of General Dynamics Corporation, San Diego, CA, USA, 1958. 40 pp.

Salzer:1958:TIE

[SL58] Herbert E. Salzer and Norman Levine. Table of integers not exceeding 1 000 000 that are not expressible as the sum of four tetrahedral numbers. Mathematical Tables and Other Aids to Computation, 12 (62):141–144, April 1958. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Spencer:1953:CPA

[Spe53] L. V. Spencer. Calculation of peaked angular distributions from Legendre polynomial expansions and an application to the multiple scattering of charged particles. Physical Review (2), 90(1):146–150, April 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Squire:1960:RM

[Squ60]  William Squire.  Relax-

ation methods. Mathematics Magazine, 33(4):177–183, March/April 1960. CODEN MAMGA8. ISSN 0025-570X (print), 1930-0980 (electronic).

Salzer:1960:TOI

[SST60] Herbert E. Salzer, Dexter C. Shoultz, and Elizabeth P. Thompson. Tables of Osculatory Integration Coefficients. Convair Astronautics, Division of General Dynamics Corporation, San Diego, CA, USA, 1960. 43 pp.

Sterne:1953:ANS

[Ste53] Theodore E. Sterne. The accuracy of numerical solutions of ordinary differential equations. Mathematical Tables and Other Aids to Computation, 7 (43):159–164, July 1953. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Stenger:1968:KPE

[Ste68] Frank Stenger. Kronecker product extensions of linear operators. SIAM Journal on Numerical Analysis, 5(2):422–435, June 1968. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stewart:1971:HRW

[Ste71] N. F. Stewart. A heuristic to reduce the wrapping effect in the numerical solution of x′ = f(t, x). BIT (Nordisk tidskrift for informationsbehandling), 11(3):328–337, September 1971. CODEN BITTEL, NBITAB. ISSN 0901-246X.

Steinberg:1972:NSV

[Ste72] J. Steinberg. Numerical solution of Volterra integral equation. Numerische Mathematik, 19(3):212–217, June 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Struble:1960:TUQ

[Str60] George Struble. Tables for use in quadrature formulas involving derivatives of the integrand. Mathematics of Computation, 14(69):8–12, January 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

links.jstor.org/sici?sici=
0025-5718(196001)14:69<8:
TFUIQF>2.0.CO;2-T&origin=
MSN.

Southwell:1953:SEP

[SV53] R. V. Southwell and Gillian Vaisey. On some eigenvalue problems of exceptional difficulty, exemplified by a case of elastic instability. Quarterly Journal of Mechanics and Applied Mathematics, 6(4):453– 480, 1953. CODEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Thompson:1959:BRN

[Tho59] G. T. Thompson. Book review: The Numerical Solution of Two-Point Boundary Problems in Ordinary Differential Equations By L. Fox. Oxford University Press, New York, 1957. xi + 371 pp. SIAM Review, 1(2):177–178, July 1959. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Takahasi:1970:EEN

[TM70] Hidetosi Takahasi and Masatake Mori. Error estimation in the numerical integration of analytic functions. Report of the Computer Centre. University of Tokyo, 3:41–108, 1970. CODEN TUCRA4. ISSN 0564-8742.

Takahasi:1981:PCG

[TN81] Hidetosi Takahasi and Takashi Nodera. Preconditioned conjugate gradient algorithms for nonsymmetric matrix. Keio Mathematical Seminar Reports, (6):51–70, 1981. ISSN 0388-

Tocher:1955:BRBc

[Toc55] K. D. Tocher. Book review:

A Short Table for Bessel Functions of Integer Orders and Large Arguments, by L. Fox; Table of Binomial Coefficients, by J. C. P. Miller. Journal of the Royal Statistical Society. Series A (General), 118 (1):106–107, ???? 1955. CO-DEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/
stable/2342530.

Todd:1959:BRF

[Tod59] John Todd. Book review:

L. Fox. The use and construction of mathematical tables. National Physical Lab., Math. Tables, v. 1, Her Majesty’s Stationery Office, London, England, 1956, 60 p., 27.5 cm. Mathematics of Computation, 13(65): 61–64, January 1959. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.
jstor.org/stable/2002204.

Todd:1964:BRN

[Tod64] John Todd. Book review: Numerical Solution of Ordinary and Partial Differential Equations edited by L. Fox, 509 pages; Addison-Wesley Publishing Co., 1962. American Scientist, 52(1):116A– 117A, March 1964. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic). URL https://www.
jstor.org/stable/27838969.

Todd:1975:CLC

[Tod75a]  John Todd.  Corrigendum:

“The Lemniscate Constants”. Communications of the ACM, 18(8):462, August 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Tod75b].

Todd:1975:LC

[Tod75b] John Todd. The lemniscate constants. Communications of the ACM, 18(1):14–19, January 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Collection of articles honoring Alston S. Householder. See corrigendum [Tod75a]. Reprinted in [Tod00].

Todd:2000:LC

[Tod00]  John Todd.  The lemnis-

cate constants. In Berggren et al. [BBB00], pages 412– 417. ISBN 0-387-98946-3 (hardcover). LCCN QA484 .P5 2000. Reprint of [Tod75b].

Treuenfels:1962:SLR

[Tre62] P. Treuenfels. Schranken für Lösungen von Randwertaufgaben. (German) [Bounds for solutions to boundary-value problems]. Numerische Mathematik, 4(1):21–23, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tsao:1975:IRT

[Tsa75] Nai-Kuan Tsao. On iterative refinement of triangular linear algebraic systems. Journal of The Franklin Institute, 299 (6):409–416, June 1975. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

Usmani:1972:MSM

[Usm72] Riaz A. Usmani. A multi-step method for the numerical integration of ordinary differential equations. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 23(3):465– 483, May 1972. CODEN ZAM-PDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Uchida:1956:RFB

[UY56] Shigeo Uchida and Michiru Yasuhara. The rotational field behind a curved shock wave calculated by the method of flux analysis. Journal of the Aeronautical Sciences, 23(9):830– 854, 1956. CODEN JASCAR. ISSN 0095-9812.

Vejvoda:1960:PBV

[Vej60]  O. Vejvoda.  Perturbed

boundary-value problems and their approximate solution. In Symposium on the numerical treatment of ordinary differential equations, integral and integro-differential equations (Rome, 1960), pages 37–41. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1960.

Viswanathan:1957:SPE

[Vis57] R. V. Viswanathan. Solution of Poisson’s equation by relaxation method—normal gradient specified on curved boundaries. Mathematical Tables and Other Aids to Computation, 11(58): 67–78, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Walsh:1966:NAI

[Wal66] Joan E. Walsh, editor. Numerical Analysis, an Introduction. Based on a symposium organized by the Institute of Mathematics and Its Applications, Birmingham, England, 1965. Academic Press, New York, USA, 1966. xiv + 212 pp. LCCN QA297 .N84 1967.

Warner:1957:SJP

[War57] F. J. Warner. On the solution of “jury” problems with many degrees of freedom. Mathematical Tables and Other Aids to Computation, 11(60):268–271, October 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Wigley:1969:MSS

[Wig69] Neil M. Wigley. On a method to subtract off a singularity at a corner for the Dirichlet or Neumann problem. Mathematics of Computation, 23(106):395–401, April 1969. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wilson:1959:SCL

[Wil59] L. B. Wilson. Solution of certain large sets of equations on Pegasus using matrix methods. The Computer Journal, 2(3):

130–133, March 1959. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wittmeyer:1958:NMD

[Wit58] H. Wittmeyer. A new method for developing simple formulae for the eigenvalues of linear ordinary self-adjoint differential equations. Journal of the Society for Industrial and Applied Mathematics, 6(2):111– 143, June 1958. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Yamamoto:1968:LAT

[Yam68]  Tetsuro Yamamoto.  On

Lanczos’ algorithm for tridiagonalization. Journal of Science of the Hiroshima University. Series A-I. Mathematics, 32:259–284, 1968. ISSN 0386-3026.

Young:1971:BRC

[YFM71] A. Young, L. Fox, and D. F.

Mayers. Book reviews: Convexity and Optimization in Finite Dimensions I, by Josef Stoer and Christoph Witzgall, 1970; On Round-off Errors in Linear Programming, by H. Müller-Merbach, 1970; Approximation Theory, edited by A. Talbot, 1970. The Computer Journal, 14(4):390, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/

                                        Volume_14/Issue_04/tiff/
                                        390.tif.

Young:1956:REA

[You56] Robert L. Young. Report on experiments in approximating the solution of a differential equation. Journal of the ACM, 3 (1):26–28, January 1956. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Young:1971:ISL

[You71] David M. Young. Iterative Solution of Large Linear Systems. Computer Science and Applied Mathematics, Editor: Werner Rheinboldt. Academic Press, New York, USA, 1971. ISBN 0-12-773050-8. xxiv + 570 pp. LCCN QA195 .Y681 1971.

Zaroodny:1955:ERM

[Zar55] S. J. Zaroodny. An Elementary Review of the Mathieu–Hill Equation of Real Variable Based on Numerical Solutions. Ballistic Research Laboratories, Aberdeen Proving Ground, MD, USA, 1955. 29 pp. Memo. Rep. No. 878.

Zadunaisky:1980:IIA

[ZL80] P. E. Zadunaisky and G. Lafferriere. On an iterative improvement of the approximate solution of some ordinary differential equations. Computers and Mathematics and Applications, 6(1):147–154, 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).